### v1

# CounterQuill: Investigating the Potential of Human-AI Collaboration in Online Counterspeech Writing

Abstract.
Online hate speech has become increasingly prevalent on social media platforms, causing harm to individuals and society. While efforts have been made to combat this issue through content moderation, the potential of user-driven counterspeech as an alternative solution remains underexplored. Existing counterspeech methods often face challenges such as fear of retaliation and skill-related barriers. To address these challenges, we introduce CounterQuill, an AI-mediated system that assists users in composing effective and empathetic counterspeech. CounterQuill provides a three-step process: (1) a learning session to help users understand hate speech and counterspeech; (2) a brainstorming session that guides users in identifying key elements of hate speech and exploring counterspeech strategies; and (3) a co-writing session that enables users to draft and refine their counterspeech with CounterQuill. We conducted a within-subjects user study with 20 participants to evaluate CounterQuill in comparison to ChatGPT. Results show that CounterQuill’s guidance and collaborative writing process provided users a stronger sense of ownership over their co-authored counterspeech. Users perceived CounterQuill as a writing partner and thus were more willing to post the co-written counterspeech online compared to the one written with ChatGPT.
Human-AI Collaboration, counterspeech, large language model, writing assistants.
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
Figure 1. The image showcases CounterQuill’s key features: (A) Learning Session (left) - Educates users about hate speech, counterspeech, and their impact through interactive lessons and quizzes. (B) Brainstorming Session (middle) - Guides users to identify key elements of hate speech, such as targeted identities and dehumanizing actions, and collaboratively brainstorm counterspeech strategies. (C) Writing Session (right) - Provides an AI-assisted writing environment with grammar checking, empathetic tone polishing, and personalized suggestions to help users write counterspeech.
1. introduction
The rapid growth of online communication platforms has led to an increase in the prevalence of hate speech, causing significant harm to individuals and society (Holt and Bossler, 2020). A study by the Pew Research Center found that 41% of U.S. adults have experienced online hate speech, with 18% subjected to severe forms such as physical threats, persistent harassment, or sexual harassment (DUGGAN, 2017). Youth and young adults are even more likely to be exposed to online hate. One cross-national study of online hate in Finland, France, Poland, Spain, the United Kingdom, and the United States found that approximately 70% of youth and young adults ages 18 to 25 had seen online hate materials. Exposure rates ranged from 65% in France to almost 79% in Finland, with 73% of Americans being exposed (Reichelmann et al., 2021). To reduce online hate, technology companies have employed human moderators and automated moderation systems. However, prior research suggests that these efforts, particularly the practice of deplatforming – banning users who express hateful views – may not effectively eliminate hate speech and often drives users to less monitored online environments (Horta Ribeiro et al., 2023). The limitations of traditional moderation methods have prompted scholars to explore the potential of user-driven counterspeech as an alternative solution to combat online hate.
Counterspeech is “communication that tries to counteract potential harm brought about by other speech” (Cepollaro et al., 2023; Rieger et al., 2018). With respect to hate speech, counterspeech is a direct response to derogatory or harmful content, intended to undermine or refute the hateful message with the intention of stopping it, reducing its impact, or providing support to those it targeted (Rieger et al., 2018). However, individuals often face significant challenges and barriers when engaging in this practice, such as fear of retaliation as well as emotional and skill-related barriers (Mun et al., 2024; Ping et al., 2024). Previous research has identified such barriers as crucial factors deterring people from engaging in counterspeech (Ping et al., 2024). Individuals with these barriers were found to be less satisfied with their own counterspeech, experienced more difficulty in writing it, and were less likely to perceive their counterspeech as effective (Ping et al., 2024).
Given the barriers that individuals face when engaging in counterspeech, the rapid advancements in Natural Language Processing (NLP) and Large Language Models (LLMs) offer avenues for supporting individuals in combating online hate (Yu et al., 2022; Zhu and Bhat, 2021; Saha et al., 2022; Garland et al., 2020). Researchers have been exploring various aspects of NLP techniques and LLMs, such as context-based hate speech recognition (Yu et al., 2022; Garland et al., 2020) and counterspeech generation (Zhu and Bhat, 2021; Saha et al., 2022). However, despite such potential, LLM-generated counterspeech have several limitations. For example, some worry that AI-generated counterspeech might lack personal touch(Baider, 2023; Ping et al., 2024; Mun et al., 2024). Prior studies show that some also doubt AI’s ability to mimic human emotions such as empathy, which is important when engaging in sensitive topics and countering hate speech (Ping et al., 2024; Hangartner et al., 2021). Furthermore, previous research discovered that counterspeech generated by AI (such as ChatGPT) prevented users from feeling a sense of ownership over the words and ideas expressed because ChatGPT automates the generation without engaging people in the process of writing counterspeech (Kumar, 2024).
To address these challenges, we introduce CounterQuill, an AI-mediated counterspeech writing partner that assists users in composing counterspeech by incorporating their own experiences and understanding. The system offers a co-writing feature that provides writing assistance based on the user’s ideas, allowing them to explore different strategies and tones, such as an empathetic approach. Throughout the interactive process, CounterQuill offers targeted feedback on each user’s counterspeech idea, refining and rewriting the text based on the user’s focus and preferences. By engaging with the system’s feedback, users discover that a user-centric approach — combining AI-provided advice with human-written personal experiences — creates a tailored and personally relevant counterspeech.
To evaluate CounterQuill, we conducted a within-subjects study with 20 participants. Each participant was shown online hate posts and tasked to write counterspeech under two different conditions: (1) using ChatGPT, and (2) using CounterQuill. To minimize potential biases, we randomly assigned the order of conditions: with half the participants starting with our system and the other half beginning with ChatGPT. Specifically, we address the following research questions in our user study to better understand the role of AI in assisting users learn, brainstorm, and draft counterspeech: (RQ1) How does the learning session influence users’ understanding of hate speech, as well as the strategies and empathy-based approaches for counterspeech? (RQ2) How does the brainstorming session guide users through the process of identifying and responding to identity-based dehumanization in online hate speech impact users’ confidence, and perceived ability to engage in counterspeech? (RQ3) How does AI assistance in drafting and refining counterspeech impact users’ perceived self-efficacy in counterspeech writing and their perspectives towards the counterspeech co-written with AI?
Contributions: The intellectual merits of this work lies in the design of an AI-mediated counterspeech system and the potential social impact of human-AI collaboration.
  • We show that CounterQuill assists users in identifying the core elements of hate speech, such as the targeted individual or group’s identity and the dehumanizing actions. The system’s highlighting practice and AI-generated feedback enhance participants’ ability to recognize these key components.
  • We find that CounterQuill increases users’ confidence in crafting empathetic counterspeech. The system’s brainstorming session combined with the co-writing session that provides targeted suggestions and rewriting assistance, empowers users to compose empathetic counter responses.
  • We uncover a relationship between user cognitive effort and behavioral outcomes in AI-mediated communication (AIMC) systems. Our findings suggest that when users invest more time and effort in collaborating with an AI system to craft counterspeech, they develop a stronger sense of ownership over the co-authored counterspeech, increasing their likelihood of posting it on social media.
  • Finally, our system demonstrates broad adaptability to diverse user groups, as it can assist different individuals in countering hate speech. By providing customized guidance and support tailored to users’ unique experiences, attitudes, and skill levels, CounterQuill showcases the potential for AI-mediated tools to help users combat online hate.
2. Related work
2.1. Online Hate and its Dangers
Online hate – or cyberhate – is a type of cyberviolence that uses computer technology to express hatred toward a group based on the group’s race, ethnicity, gender, gender identity, sexual orientation, national origin, religion, or other group characteristic (Perry, 2009; Räsänen et al., 2016). It differs from cyberbullying because of this collective focus. Like terrorism, online hate can advocate the hatred of one group or the hatred of numerous groups. Exposure to online hate has increased in the U.S. (Costello and Hawdon, 2020), with over 60 percent of Americans ages 18 to 25 being exposed to it in 2023 (Costello et al., 2023; Hawdon, 2024). But it is not limited to the U.S., it is a global phenomenon (Reichelmann et al., 2021; Commission, 2018). Although not everyone who sees online hate suffers negative consequences, exposure to it is correlated with both short-term and long-term emotional distress, such as mood swings, anger, feelings of isolation, decreased levels of trust, and fear (Costello and Hawdon, 2020; Foxman and Wolf, 2013; Näsi et al., 2015; Nyman and Provozin, 2019). Because they are still forming their identities, young people are especially vulnerable to negative effects of online hate (Billett, 2014). Given the amount of online hate that can be found online and the dangers it can pose, it is imperative that we find effective tools for detecting it and combatting it. Recent advances in automated detection have been made, but detecting it is not enough. We also need to find ways to lessen its impact. Combatting online hate with counterspeech is one such strategy.
2.2. Human-AI Interaction with Generative AI
Recent advancements in Generative AI like GPT-4, Bard, and Llama have enabled their integration into various application domains (Achiam et al., 2023; Rahsepar et al., 2023; Touvron et al., 2023), causing interest in how users interact, collaborate, and co-write with these AI systems (Wu et al., 2023; Tang et al., 2024; Wang et al., 2023; Duin et al., 2023; Balel, 2023; Lee et al., 2024a, 2022). In the field of co-writing with AI, researchers have shed light on aspects of user engagement, adaptation, and reliance on AI-generated content (Wu et al., 2023; Lewis and Moorkens, 2020; Lee et al., 2024a). These studies clarify activities ranging from using AI suggestions for conceptualization and content generation to evaluating (Desmond et al., 2024; Lee et al., 2024b), editing, and refining machine-generated text (Yang et al., 2022; Balel, 2023; Coenen et al., 2021; Zhang et al., 2023), such as creative writing (Coenen et al., 2021; Yang et al., 2022), technical documentation (Zhang et al., 2023; Duin et al., 2023), and online communication (Jakesch et al., 2023). Despite such advancements, there are limitations in AI-assisted writing, such as incorporating user instructions (Yeh et al., 2024; Wasi et al., 2024), handling biased AI-generated content (Jakesch et al., 2023), and navigating challenges related to AI’s understanding of user intent or desired tone (Ping et al., 2024; Lee et al., 2024a). For instance, Osone et al. (2021) identified challenges in collaborating with AI in creative writing, where experienced writers exhibited dissatisfaction with AI-generated content due to misalignment with expected storylines (Osone et al., 2021). Meanwhile, as Zhou et al. (2023) revealed that generative AI creates disinformation by both reshuffling human-made data and manipulating language to produce content that seems credible to targeted audiences.(Zhou et al., 2023).
As the integration of AI into writing platforms becomes more prevalent, not only writers and researchers but also social media developers and users will incorporate AI tools into their workflows to enhance the co-creation of social media content (Mun et al., 2024; Ping et al., 2024). Particularly, the collaboration between humans and AI in generating counterspeech involves a nuanced understanding of language, tone, and context to counteract harmful narratives online (Mun et al., 2024; Ping et al., 2024; Ray, 2023). For instance, Ray (2023) and Ping et al. (2024) highlighted that researchers, while crafting counterspeech, find AI tools beneficial in generating initial ideas. Still, the refinement of language, ensuring accuracy, and constructive tone often require human intervention (Ping et al., 2024; Ray, 2023). Additionally , the design and implementation of AI-assisted tools for counterspeech present challenges, such as ensuring that the AI-generated content involves sufficient human participation (Ping et al., 2024; Ray, 2023), aiding users in providing ample background and support when co-writing counterspeech with AI (Ping et al., 2024), and assisting users in maintaining the credibility and efficacy of the counter-argument when crafting counterspeech (Mun et al., 2024; Ping et al., 2024).
Overall, while previous research has provided valuable support for initial counterspeech creation, conscious human engagement is needed to ensure its accuracy and effectiveness in the Human&AI co-creation process (Mun et al., 2024; Ping et al., 2024). Our research aims to bridge the gap by exploring ways to enhance human-AI collaboration so that AI-generated counterspeech resonates effectively with the target audience and context.
2.3. Counterspeech Generation
Text generation, especially counterspeech generation, is one of the many open challenges in NLP research that has made breakthroughs in recent years(Zhu and Bhat, 2021; Chung et al., 2020, 2021; Saha et al., 2024). Researchers have explored various aspects of counterspeech generation using LLMs, such as generating contextually relevant responses (Hassan and Alikhani, 2023; Zhu and Bhat, 2021), knowledge-grounded counterspeech generation (Chung et al., 2021), and ensuring that the generated counterspeech adheres to ethical and societal norms (Mun et al., 2023). For instance, Chung et al. (2022) explored whether adding relevant external information could improve LLMs’ ability to generate counterspeech to hate speech (Chung et al., 2021). By extracting keyphrases from the hate speech and finding related outside knowledge, they were able to generate more suitable and informative counterspeech. Meanwhile, Hassan et al. (2023) have developed a novel framework based on theories of discourse to study the inferential links connecting counterspeech to hateful comments (Hassan and Alikhani, 2023). They proposed a taxonomy of counterspeech derived from discourse frameworks and discourse-informed prompting strategies to generate contextually-grounded counterspeech (Hassan and Alikhani, 2023).
Despite LLMs’ ability to generate counterspeech, we do not know whether machine-generated counterspeech aligns with people’s preferences or whether people would actually use it to counter hate online (Mun et al., 2024; Ping et al., 2024). Prior research shows that people are hesitant to use counterspeech entirely generated by AI because it strips away their personal voice and raises concerns about authenticity (Ping et al., 2024). On the other hand, some believe AI assistance in counterspeech-writing can reduce the emotional burden of responding to online hate and help them articulate their thoughts better (Ping et al., 2024). Researchers in HCI and CSS emphasize the importance of a human-centered approach in understanding the use of AI-generated content. Likewise, we argue that effectively using LLMs to counter hate speech in the real world requires attention to human involvement, preferences, and concerns. Our work adopts this approach to improve LLMs’ ability to generate human-centered counterspeech and to explore how humans can co-create counterspeech with AI in alignment with their preferences, personal voices, and ideas.
2.4. Ownership in Human-AI Collaboration
Previous research has demonstrated that ownership extends beyond physical or tangible objects, encompassing everything from people to text to ideas (Heider, 1958; Mieczkowski, 2022). Regarding how interaction with AI systems affects the human sense of ownership, past studies have found that the more people engage in their interactions with AI, the more likely they are to develop a strong sense of ownership over the outcomes of their co-creation with the AI (Louie et al., 2021; Pierce et al., 2003; Louie et al., 2020a). For instance, in Louie et al.’s (2021) study of AI music co-creation, participants indicated that they felt the final product did not belong to them due to their minimal control over the AI-generated suggestions (Louie et al., 2021). Furthermore, Draxler et al.’s study of AI-assisted writing suggested that users’ perceived control over AI-generated content and their feeling of leadership in the interaction significantly impact their sense of ownership (Draxler et al., 2024).
As AI-assisted writing continues to evolve, people are increasingly using with AI to create content on social media; one such example is counterspeech writing (Buerger, 2021; Saha et al., 2022; Mun et al., 2024). While previous research has shown that AI can help people overcome some writing barriers in the process of creating counterspeech (Mun et al., 2024; Zhu and Bhat, 2021; Saha et al., 2024), studies by Kumar (2024) and Ping et al. (2024) have highlighted that people feel AI, such as ChatGPT, prevented users from feeling a sense of ownership over the words and ideas expressed (Kumar, 2024; Ping et al., 2024). This loss of ownership has also raised concerns about authenticity and ethics (Kumar, 2024), which may limit their willingness to use the generated counterspeech.
Overall, the growing collaboration between humans and AI in content creation has led to an interplay of user effort (Mieczkowski, 2022; Louie et al., 2021), perceived ownership (Mieczkowski, 2022; Kumar, 2024; Louie et al., 2020a; Pierce et al., 2003), and ethical considerations (Kumar, 2024; Ping et al., 2024). Our research aims to further explore the relationship between people’s effort, perceived ownership, and future behavioral outcomes in human-AI collaboration.
3. BACKGROUND AND DESIGN GOALS
We take a human-centered approach in the design of CounterQuill as an interactive human-AI system that empowers users with relevant knowledge (DG1), ideation space (DG2), and collaborative writing assistance (DG3) rather than simply generating counterspeech. By doing so, our system aims to help users engage in the process of crafting effective and empathetic counterspeech that incorporates their preferences and thought-process in countering online hate, rather than soleley delivering a machine-generated output. Our system is designed to achieve the following design goals:
  • DG1. Help users understand hate speech and counterspeech and their impact.
  • DG2. Help users ideate counterspeech ideas through a brainstorming process.
  • DG3. Help users draft and refine their counterspeech with AI assistance.
3.1. Help Users Learn About Hate Speech & Counterspeech and Their Impact (DG1)
3.1.1. Understanding Hate Speech and its Impact
Previous research found that many internet users, especially younger generations, may not fully understand what constitutes hate speech or its harmful effects (Räsänen et al., 2016; Anti-Defamation League, 2019). A survey conducted by the Anti-Defamation League revealed that 65% of American youth were exposed to online hate speech, but only 49% of them recognized and reported hateful content to social media companies (Anti-Defamation League, 2019). This lack of awareness underscores the need for educational initiatives to help users identify hate speech effectively. To address this issue, NGOs have launched campaigns aimed at raising awareness about hate speech and teaching people how to recognize it (Laville, 2016; Schieb and Preuss, 2016).
Previous research has found that hate speech can be expressed both explicitly and implicitly (ElSherief et al., 2021; Gunturi et al., 2024). Explicit hate speech directly attacks social groups or individuals through overt language, such as profanity (Reichelmann et al., 2021; Howard, 2021). In contrast, implicit hate speech is more subtle (ElSherief et al., 2021), often relying on coded language (ElSherief et al., 2021; Ocampo et al., 2023), neologisms (Retta, 2023), sarcasm (Ocampo et al., 2023), or microaggressions (Gunturi et al., 2024) to convey prejudice or discrimination. Thus, same words or phrases can have different meanings depending on the specific situation, the speaker’s intent, and the audience’s perception (Gunturi et al., 2024). For such reasons, it can be particularly difficult for humans to recognize implicit hate speech.
3.1.2. Understanding counterspeech and its Impact
Furthermore, previous research found that the lack of clear guidelines and practices for crafting effective counterspeech is another factor contributing to people’s reluctance to respond to hate speech (Ping et al., 2024). Baider (2020) observed that individuals often struggle to determine what constitutes impactful counterspeech and how to formulate responses that can make a difference (Baider, 2020).
Research has highlighted the potential for unintended consequences when counterspeech is not delivered carefully (Howard, 2021). As Howard’s (2019) study found, counterspeech perceived as aggressive or preachy can sometimes backfire, leading to increased hate speech (Howard, 2021). Humorous counterspeech also has mixed results on hate speech. While humour can sometimes defuse hate speech, attempts at humorous counterspeech in other cases can backfire (Schieb and Preuss, 2016). Compared to other strategies like humour or warning of consequences, Hangartner et al. show that empathy-based counterspeech - counterspeech that humanizes the victim and reminds the perpetrator that people can be hurt by their behaviour - is significantly more likely to lead to apologies or deletion of hateful posts by bad actors (Hangartner et al., 2021). However, most people are not only unaware of the different counterspeech strategies and their potential pitfalls and risks, but which ones may be more effective (Ping et al., 2024). To address these issues, our system is designed to help people learn more about hate speech, counterspeech and their impact.
3.2. Help Users Ideate Counterspeech Ideas Through a Brainstorming Process (DG2)
Previous research has found that while many people recognize the harmful effects of hate speech, they may feel uninformed or unable to engage in counterspeech effectively. Studies by Mun et al. (2024), Buerger et al. (2021), and Ping et al. (2024) revealed that due to cognitive load, users often hesitate or struggle to create effective counterspeech (Mun et al., 2024; Buerger, 2021; Ping et al., 2024) . This cognitive load arises from the difficulty in accurately identifying the key hate elements in hate speech, such as aspects of an individual or group’s identity and actions or perceptions that dehumanize that person or group (Waltman, 2018; Harel et al., 2020). Furthermore, many people may lack the skills to effectively counter hate speech (Ping et al., 2024). A study by Costello et al. (2017) found that although young adults (from an online survey, N = 963) recognized the importance of countering online hate speech, they were often uncertain about how to effectively counter the key hate elements of hate speech (Costello et al., 2017).
To address these challenges, we adopted the traditional brainstorming approach (TBS) (Gallupe et al., 1991; Yu-Han and Chun-Ching, 2023) and extended it by drawing on prior HCI work on AI technologies that facilitate brainstorming (Lavrič and Škraba, 2023; Zhang et al., 2023; Yu-Han and Chun-Ching, 2023; Cepolina et al., 2022). Researchers in these studies use LLMs to help people generate (Zhang et al., 2023; Yu-Han and Chun-Ching, 2023), refine (Yu-Han and Chun-Ching, 2023; Geerts et al., 2021), and evaluate ideas (Cepolina et al., 2022). For example, Gero et al. developed Metaphoria, an LLM-powered system that assists writers in brainstorming creative ideas by generating metaphorical associations and encouraging diverse creative outputs (Gero et al., 2022). In addition, Zhang et al. created Visar, a human-AI collaborative writing assistant that helps writers brainstorm and revise hierarchical goals within their argumentative writing context (Zhang et al., 2023).
Lubart (2005) and Chiu et al. (2023) (Lubart, 2005; Yu-Han and Chun-Ching, 2023), identified four key components in HCI brainstorming: (1) AI engaging in communication and guidance with humans during the thinking process; (2) AI providing targeted feedback; (3) AI and humans collaborating to organize brainstorming results; and (4) AI encouraging reflective behavior by integrating human-machine collaboration into the creative process. Built on previous research, we developed two AI agents: [Uncaptioned image] Brainstorming Guidance Agent (BGA), which facilitates communication, guidance, and encouragement, and [Uncaptioned image] Brainstorming Feedback Agent (BFA), which provides feedback and collaborates with users to organize the outputs of the brainstorming process. Overall, our brainstorming session aims to help users ideate counterspeech ideas through a brainstorming process assisted by these AI agents.
3.3. Help Users Draft and Refine Their Counterspeech With AI Assistance (DG3)
Previous research has highlighted the difficulties of writing effective counterspeech (Ping et al., 2024; Howard, 2021). Studies by Mathew et al. found that a significant portion of counterspeech on Twitter were ineffective or even counterproductive, sometimes escalating conflict or using inappropriate language (Mathew et al., 2020, 2018). Ping et al. (2024) reveal the various challenges individuals face when writing online counterspeech, such as the emotional burden, time commitment, and perceived lack of skill in crafting effective responses (Ping et al., 2024). Their findings showcase that these barriers are especially pronounced for women, who report greater uncertainty in their ability to write impactful counterspeech and experience a higher emotional toll in the process (Ping et al., 2024).
Furthermore, Hangartner (2021) emphasized the importance of empathy in combating online hate speech, finding that empathy-based counterspeech consistently reduced hate speech compared to other strategies like humour or warning of consequences (Hangartner et al., 2021). This finding underscores the importance of empathy in combating online hate speech, suggesting that counterspeech that encourages empathy can reduce hostility towards marginalized groups (Hangartner et al., 2021).
To address these challenges and provide guidance on empathetic counterspeech, we developed two AI agents,[Uncaptioned image] Writing Guidance Agent (WGA) and [Uncaptioned image] Writing Assistance Agent (WAA), for an AI-human co-writing session to increase user confidence and self-efficacy while reducing the emotion-related burden of writing counterspeech. The session also provides users with rewriting features to refine their writing, such as adjusting the tone to be more empathetic.
4. System Implementation
CounterQuill was developed using React for the front end, and Python + MySQL for the back end, which was served through Flask. The system’s interface was designed with a minimal aesthetic using the Tailwind CSS framework to create clear indicators of system features.
4.1. CounterQuill Learning Session
Figure 2. User Interface for DG1 (Help people understand hate speech & counter speech and their impact). (1) Section for educating people to understand hatespeech; (2) Section for educating people to understand counterspeech; (3) Test people’s understanding of hatespeech & counterspeech.
Figure 3. User Interface for educating people to understand hate speech: (1) The definition of hate speech; (2) The explicit or implicit hate speech, and (3) How does hate speech affect us?
As shown in Figure 2, the Learning session has three main sections: 1 What is Hate Speech?
2 Learning about Counterspeech and 3 Testing your Understanding. On the left side of the accordions is our CounterQuill guidelines message, which summarizes the main tasks during the Learning session as well as the basic steps to follow. We designed the interface of Learning session using classic CSS accordions, a common user interface pattern that allows users to interactively show and hide sections of content on a web page. The content for the Learning session was developed through an iterative design process informed by literature and expertise from the first author, two sociologists, and one HCI expert.
1What is Hate Speech?
contains three learning content subsections, which users can navigate through by clicking the Next button to progress and the Previous button to review. It covers three crucial topics for comprehending hate speech: (1) The definition of hate speech; (2) The explicit or implicit hate speech, and (3) How does hate speech affect us? (as shown in the Figure 3).
Similarly, 2 Learning about Counterspeech also features three learning content subsections, navigable via the Next and Previous buttons. It focuses on three key aspects of counterspeech: (1) What is counterspeech? (2) What makes counterspeech effective? and (3) What is an empathy-based counterspeech? (as shown in the Figure 4).
Figure 4. User Interface for learning about counterspeech: (1) What is counterspeech? (2) What makes counterspeech effective? and (3) What is an empathy-based counterspeech?
Lastly, 3 Testing your Understanding includes four questions to assess users’ comprehension of the learning material covered in the previous sections. These questions evaluate users’ understanding of (1) the definition of hate speech, (2) the effects of explicit and implicit hate speech, (3) the definition of counterspeech, and (4) the concept of empathy-based counterspeech (see Appendix C for detailed quiz questions).
4.2. CounterQuill Brainstorming Session
To help users ideate counterspeech ideas, we designed the Brainstorming session with two key sections: (a) a section for users to highlight key elements of hate speech, as shown in Figure 5, and (b) a section to encourage users to engage in thinking about hate speech by considering the negative assumptions suggested by the statement and the feelings of those targeted (Figure 6). The brainstorming session includes: 1 Tutorial on highlighting practice, 2 Highlighing area, 3 Brainstorming through Q&A and 4 CounterQuill suggestions on brainstorming results.
Figure 5. User Interface for DG2 (Help users ideate counterspeech ideas through a brainstorming process)’s Highlighting Practice: (1) Tutorial on highlighting practice; (2) Highlighting area: identify the individual/ group’s identity and dehumanizing actions.
4.2.1. Highlighting Practice
For 1 , the tutorial features a [Uncaptioned image] BGA that communicates the core brainstorming tasks and steps. It dynamically demonstrates how to highlight aspects of an individual or group’s identity in yellow and actions that dehumanize them in green. The tutorial, developed by the first author, two sociologists and an HCI scholar, uses a sample hate speech statement (“I saw a random black man jogging nearby our house today and now I feel unsafe walking around my own neighborhood”).
Figure 6. User Interface for DG2’s Question Answering: (1) Brainstorming through Q&A, (2) CounterQuill suggestion on brainstorming results.
In 2 Highlighing area, users manually select and highlight the identity aspects in yellow and the dehumanizing actions in green. Once users have completed highlighting the content, they can click the ”Done” button, triggering [Uncaptioned image] BFA to semantically compare the user-highlighted content with the system’s predefined correct answer (Luo et al., 2023). BFA captures the user-highlighted text and sends it to the backend service, where it is compared using the GPT-3.5-turbo model with the following prompt:
{ ”role”: ”assistant”,
”content”:
      ”context”: ”The correct individual or a group’s identity and dehumanizing action in the speech
       text ‘{hatespeech}’ is: ‘{identity}’ and ‘{action}’ ”,
      ”query”: If the user selects ‘{user selection_1}’ as identity and ‘{user selection_2}’ as dehumanizing
       action, are they semantically equivalent to the correct answer?”,
      ”responseOptions”: [”Yes”, ”No”] }
For example, in the hate speech statement “I saw a random black man jogging nearby our house today and now I feel unsafe walking around my own neighborhood”, the correct semantic understanding of the dehumanizing action could be “feel unsafe”. If the user selects “unsafe” or “I feel unsafe”, they have correctly understood the dehumanizing action, and the response based on this prompt will be ‘Yes’. However, if the user selects “jogging nearby”, it is an incorrect understanding of the dehumanizing action, and the response based on this prompt will be ‘No’ and the BFA will generate corresponding corrective feedback: CounterQuill highlighting feedback. Additionally, users can click the “View” button to see the differences between their understanding of identity and dehumanizing action compared to the system’s interpretation.
4.2.2. Question Answering
For 3 Brainstorming through Q&A, the first author collaborated with two sociologists and an HCI scholar to design two thought-provoking questions that encourage users to provide thoughtful written responses (Figure 6). Question 1 asks, “What negative stereotypes or assumptions about the targeted group or individual are suggested by the statement?” This question aims to let users think of harmful assumptions implied about the targeted group or individual. Question 2 asks, “Consider the feelings and experiences of someone who identifies with the group mentioned in the statement. How might this comment affect their sense of safety, belonging, or self-esteem?” This question encourages users to reflect on the impact of identity-based dehumanization. Users can input their responses in the provided text box. For example, in Figure 6, the user has written, “The statement implies harmful stereotypes: the presence of black people in a neighborhood leads to decreased property values and a decline in the overall quality of the area”.
For 4 CounterQuill suggestions on brainstorming results, after completing their input for both questions, users can click the “Done” button, prompting CounterQuill to provide suggestions based on the user’s responses. These suggestions include (1) feedback on the user’s understanding and (2) suggestions on strategies for crafting their counterspeech. To implement this step, clicking the ”Done” button triggers the BFA, which captures the user’s input text and sends it to the backend service. The backend service then uses the GPT-3.5-turbo model with the following prompt:
{ ”role”: ”assistant”,
”content”:
       ”task”: “Your task is to generate thoughtful and constructive counterspeech suggestions in response
       to a user’s answer to the question: ‘{QUESTION}’
             - Begin by respectfully acknowledging the user’s contribution and perspective. If the user
            expresses offensive or harmful views, kindly explain why such statements are problematic.
- Next, provide a concise and insightful evaluation of the user’s response. Offer either a
            thought-provoking observation that encourages further reflection or constructive feedback
            that helps the user refine their argument.
- Conclude by offering clear and actionable advice on crafting effective counterspeech related
            to the issue at hand. ” }
After the GPT-3.5-turbo model processes the user’s input, the generated suggestions are displayed on the frontend. Users can then read CounterQuill’s suggestions and select the parts they find noteworthy and helpful for writing their counterspeech. By selecting the desired content and clicking the “Take Notes” button, users can highlight the text to create notes. To enable users to highlight selected data as notes, clicking the “Take Notes” button triggers the BFA to capture the user-selected content and store it in the backend database. Once stored, the user-selected content (data) is automatically highlighted in blue, as shown in (Figure 6).
4.3. CounterQuill Co-Writing Session
To help users draft and refine their counterspeech, we designed a co-writing session with CounterQuill. This session includes four main sections: 1 Tutorial about CounterQuill writing assistant, 2 Brainstorming notes section for users to review, 3 Text area for users to write counterspeech, and 4 CounterQuill writing assistance tool.
In section 1 , the [Uncaptioned image] WGA interacts with users to introduce the writing tasks they need to complete during the session, including the hate speech they need to counter and a video tutorial on co-writing with CounterQuill. The tutorial video demonstrates how users can obtain writing suggestions and refine their counterspeech with the help of CounterQuill.
2Brainstorming notes section
displays the notes that users have taken by clicking the “Take Notes” button during the brainstorming session. These notes are the ones users find noteworthy and helpful for writing their counterspeech. To implement this feature, the WGA accesses our backend database, where user-selected content is stored (see §3.2.2). After indexing the user’s notes, the WGA sends them back to the frontend and displays them to users, allowing users to make initial self-revisions based on their notes.
Figure 7. User Interface for DG3 (Help users draft and refine their counterspeech with AI assistance). (1) Tutorial about CounterQuill writing assistan, (2) Brainstorming notes taken by the users, (3) Text area for users to write counterspeech.
In 3 Counterspeech writing area, the WGA first guides users, indicating that this section is for writing their counterspeech. The WGA then uses the user’s reflections on hate speech from the brainstorming session (i.e., their responses to the two questions considering the negative assumptions suggested by the statement and the feelings of those targeted) as an initial draft for their counterspeech. This reduces the user’s psychological pressure when drafting counterspeech from scratch. Users can then make initial additions, deletions, and modifications to the content based on their brainstorming notes. As users begin inputting content, our system automatically saves the input to the backend database. Once users complete their initial draft, they can select a portion or all of the text and click on the writing assistant button to invoke the help of the CounterQuill Writing Assistant (Figure 8 B ).
Figure 8. User Interface for DG3 - Select text and use CounterQuill writing assistant to refine the content. (A) CounterQuill Writing Help Button, (B) CounterQuill writing assistant’s features, (C) Sample result polished by CounterQuill
For 4 CounterQuill writing assistant, we have designed features for grammar checking, polishing content to an empathetic tone, and rewriting content using the user’s brainstorming notes 1 and 2. Finally, to help users customize their writing, we have designed a “Customize Your Writing” area where users can input their desired final writing goals to align with their preferences. To implement the CounterQuill writing assistant, we have prompted the GPT-3.5-Turbo model. When users click on the corresponding button (Figure 8 A ), the [Uncaptioned image] WAA sends the request represented by the button to the backend, and the GPT-3.5-Turbo model rewrites the text according to the user’s needs (see Appendix D for detailed prompts). Once CounterQuill completes the rewriting, the WAA returns the result to the frontend and presents it in (Figure 8 C ). Finally, if users are satisfied with the system result, they can click the “Insert” button to insert it into the text editing area. If users are not satisfied, they can click the “Retry” button, and the system will perform another rewrite.
5. User Study
We conducted a user study with 20 participants to evaluate CounterQuill. The study was approved by an Institutional Review Board (IRB), and designed to take 70 minutes for each participant. Our user study examined the following research questions:
  • RQ1.
    How does the learning session influence users’ understanding of hate speech and the strategies and empathy-based approaches for counterspeech?
  • RQ2.
    How does the brainstorming session guide users through the process of identifying and responding to identity-based dehumanization in online hate speech impact users’ confidence, and perceived ability to engage in counterspeech?
  • RQ3.
    How does AI assistance in drafting and refining counterspeech impact users’ perceived self-efficacy in counterspeech writing and their perspectives towards the counterspeech co-written with AI?
Table 1. Demographic information of CounterQuill user study participants.
ID Gender Ethnicity Age Education Sexual Orientation
P1 Male Asian/Asian American 21 Bachelor’s degree Heterosexual
P2 Male Black/African American 26 Some college Heterosexual
P3 Female Black/African American 24 Associate degree Heterosexual
P4 Female White/European American 32 Associate degree Heterosexual
P5 Female White/European American 44 Bachelor’s degree Bisexual
P6 Female Asian/Asian American 32 Bachelor’s degree Heterosexual
P7 Male Asian/Asian American 23 Bachelor’s degree Heterosexual
P8 Male Hispanic/Latina/o/x 21 Bachelor’s degree Heterosexual
P9 Male Asian/Asian American 23 Bachelor’s degree Heterosexual
P10 Male Middle Eastern/North African 24 Doctorate degree Bisexual
P11 Female White/European American 24 Master degree Heterosexual
P12 Female White/European American 26 Master degree Heterosexual
P13 Male Black/African American 34 Bachelor’s degree Heterosexual
P14 Female Black/African American 29 Bachelor’s degree Heterosexual
P15 Female Asian/Asian American 27 Master degree Homosexual
P16 Male White/European American 32 Master degree Heterosexual
P17 Male Middle Eastern/North African 21 Bachelor’s degree Heterosexual
P18 Male White/European American 22 Bachelor’s degree Heterosexual
P19 Male White/European American 31 Bachelor’s degree Heterosexual
P20 Female Black/African American 34 Associate degree Heterosexual
5.1. Participants
We recruited 20 native English-speaking participants (9F11M), whose ages ranged from 20 to 44 (mean = 26, SD = 5.7), through word-of-mouth, LinkedIn, Reddit, and X (Twitter). Five of them participated in the study in person at a usability lab and fifteen participated virtually through Zoom. Participants were required to complete a pre-screening survey to test whether they were required to attend our user study (See Appendix A) and collect their basic demographic information, including age, gender, sexual orientation, and race/ethnicity. We established two key criteria for participation: first, participants needed to have prior experience with generative AI tools, such as ChatGPT, Google Gemini, or Claude, though none had used CounterQuill before the study. Second, participants were required to have encountered hate speech online and have previously engaged in counterspeech. Additionally, we tried to diversify the participant group as much as possible. The demographic information of our 20 participants is shown in Table 1. Participants were informed that their responses to surveys, as well as video and audio from the interview may be used for data analysis. After the study, participants were compensated \$20 in gift card credit for their time.
Regarding their experience with AI-supported writing tools or GenAI tools, all 20 participants had prior experience with AI-supported writing tools, such as ChatGPT and Claude. Every participant was familiar with conversational interactions using generative AI tools, with 18 participants having used ChatGPT for general writing assistance, and 6 participants specifically using AI tools, like ChatGPT, for counterspeech writing.
5.2. Study Protocol
Each study session lasted approximately 70 minutes and employed a within-subjects design. Before the study, participants received a brief introductory email about the system and a tutorial video that provided a comprehensive guide on using CounterQuill’s features. During the study, participants experienced two conditions: a 10-minute period of writing counterspeech with ChatGPT, and a 50-minute period of co-writing counterspeech with CounterQuill. The detailed study procedure is described in §5.3.
Figure 9. A diagram outlining the study protocol used. The intervention ordering was counterbalanced between participants and participants engaged with different conditions. In addition, the times labelled in this figure are the average time spent by the participants.
5.2.1. Conditions
In our user study, we implemented two distinct conditions to evaluate the effectiveness of AI-assisted counterspeech writing, drawing on previous research that used GPT-based systems (Zhang et al., 2023; Reza et al., 2024; Lin et al., 2024) such as GPT-3.5 as baseline:
  • ChatGPT (Baseline): Participants used ChatGPT to write counterspeech for the hateful content. ChatGPT allowed users to input prompts and receive generated responses from the GPT-3.5-turbo language model, which is the same model used in CounterQuill.
  • CounterQuill: Participants used the full version of CounterQuill, which includes a learning session (help users understand hatespeech and counterspeech), a brainstorming session (step-by-step counterspeech ideation support), and a writing session (co-writing with CounterQuill for counterspeech).
To minimize potential biases, we used a Latin Square experimental design (Ryan and Morgan, 2007) to ensure a balanced sequence of task topics and conditions for each study session: with half the participants starting with our system and the other half beginning with ChatGPT. To assess the potential impact of the system usage order on participants’ perceptions and the overall study outcomes, we conducted Welch’s Two-Sample t-tests that accounts for unequal variances between groups (Sakai, 2016), as detailed in Tables A2 and A3. The results confirmed that the sequence—whether participants began with ChatGPT or CounterQuill—did not significantly influence their performance or feedback on the system.
Although the primary comparison focuses on the writing session, we recognize that the preceding learning and brainstorming sessions could influence the results, and we have addressed this in the limitations and future work section.
5.3. Study Procedure
At the beginning of each study session, the experimenters collected informed consent from the participants. For the user study, each participant was presented with 20 hate speech examples and asked to select one of them. The hate speech data was sourced from Ping’s (2024) research (Ping et al., 2024), which included a total of 400 hate speech instances evenly distributed across five hate speech themes: race, gender, sexual orientation, disability, and religion. To ensure a balanced representation of themes, each participant was randomly assigned 20 hate speech examples, with an average of 4 examples per theme. Table A1 (Appendix F) lists the hate speech instances used by each of the 20 participants.
The study coordinators introduced participants to the ChatGPT counterspeech writing session and provided a 3-minute tutorial on how to use CounterQuill, followed by an introduction to a CounterQuill co-writing session. Each participant completed the two writing sessions, with the order counterbalanced between participants to minimize potential biases (Greenwald, 1976; Charness et al., 2012).
In the ChatGPT (Baseline) condition, participants used ChatGPT to write counterspeech for the hateful content. Then they completed a post-study questionnaire including NASA-TLX questions and customized questions (see Appendix B for post-study questionnaire). In the CounterQuill condition, participants went through a learning session to understand hatespeech and counterspeech, a brainstorming session for step-by-step ideation support, and used the CounterQuill co-writing assistant. They then completed a post-study questionnaire. The study ended with a 20-minute semi-structured interview, during which the experimenter asked questions to understand participants’ experiences, feedback on CounterQuill features, and reflections on the two conditions. The purpose of the semi-structured interview was to explore participants’ overall experiences, elicit their specific feedback on CounterQuill and ChatGPT for counterspeech writing, and gather suggestions for improvement.
6. Result
6.1. Quantitative Analysis
6.1.1. Statistical Analysis of ChatGPT and CounterQuill with NASA TLX
Table 2. Comparison of Average Time Spent on Counterspeech Composition
Task Average Time (min)
CounterQuill Learning 8.72 $\pm$ 1.02
CounterQuill Brainstorming 10.32 $\pm$ 1.55
CounterQuill Writing 10.48 $\pm$ 1.71
ChatGPT Writing 9.11 $\pm$ 1.06
Table 3 summarizes the ratings for each dimension of the NASA Task Load Index (TLX) across two conditions: using ChatGPT (Condition A) and using CounterQuill (Condition B). We evaluated the differences between these conditions on each dimension using paired-sample $t$-tests (Mowery, 2011). To further contextualize these findings, Table 2 provides a comparison of the average time spent on counterspeech composition in each condition.
Table 3. Comparison of NASA-TLX scores between ChatGPT and CounterQuill for counterspeech writing tasks. The results of paired $t$-test show statistically significant differences in five out of six TLX subscales, with CounterQuill scoring higher in Physical Demand, Temporal Demand, Performance, and Effort, while ChatGPT scored higher in Frustration.
[…]
NASA-TLX Items $Mean\pm SD$ $t$-value $p$-value 95% CI
ChatGPT CounterQuill
Mental Demand 4.53$\pm$1.70 4.94$\pm$1.39 -1.237 0.234 [-1.117, 0.294]
Physical Demand 2.06$\pm$1.52 3.12$\pm$1.73 -3.816 0.002** [-1.647, -0.471]
Temporal Demand 2.65$\pm$1.32 4.47$\pm$1.46 -7.410 ¡0.001*** [-2.345, -1.302]
Performance 3.88$\pm$1.17 6.35$\pm$0.70 -7.177 ¡0.001*** [-3.200, -1.741]
Effort 2.53$\pm$1.07 3.82$\pm$1.67 -3.235 0.005** [-2.142, -0.446]
Frustration 3.06$\pm$1.95 1.47$\pm$0.72 3.628 0.002** [0.660, 2.516]
The analysis of the NASA-TLX dimensions revealed several notable differences between Condition A (ChatGPT) and Condition B (CounterQuill). While Mental Demand showed a slight, non-significant increase from Condition A to B, Physical Demand, Temporal Demand, and Effort all displayed statistically significant increases in Condition B. Performance also significantly improved in Condition B, with a large effect size. Conversely, Frustration was the only dimension that showed a significant decrease in Condition B compared to A.
Figure 10. The violin plot of NASA-TLX score distribution: ChatGPT vs. CounterQull.
The violin plot (Figure 10) visually represents the score distributions across the two conditions, further emphasizing the differences highlighted by the statistical analysis. The plot shows the shifts in the distribution of scores for each dimension, with higher scores for Physical Demand, Temporal Demand, Performance, and Effort in using CounterQuill, and lower scores for Frustration.
In summary, using CounterQuill requires more effort, time, and physical energy compared to using ChatGPT. However, it also leads to significantly better perceived performance outcomes, with participants reporting a 63.66% higher rate with CounterQuill (from NASA TLX ‘Performance’). This higher performance perception is not only due to their overall satisfaction with the counterspeech they co-writed but also their strong performance in the quiz after the learning session. Out of 20 participants, 15 answered all quiz questions correctly, while the remaining five answered 2, 2, 3, 3, and 3 questions correctly, resulting in an average accuracy rate of 91.25%. Additionally, CounterQuill also contributed to a 51.96% reduction in participants’ feelings of frustration compared to ChatGPT (from ‘Frustration’).
6.1.2. Quantitatively Analyze User Perceptions and Experiences with ChatGPT and CounterQuill
Our study further investigated participants’ experiences and perceptions when using ChatGPT and CounterQuill for counterspeech writing. To gain deeper insights, we designed six customized questions (see Appendix B for detailed post-study questionnaire) that focused on specific aspects of the participants’ experience, such as confidence in identifying hate speech, effectiveness of the tools in providing guidance, and satisfaction with the final counter-speech produced. The analysis of the six customized questions revealed several notable findings (Table 4):
Table 4. Comparison of participant perceptions between ChatGPT and CounterQuill for counterspeech writing assistance based on six customized questions. The results of paried $t$-test show statistically significant differences favoring CounterQuill in all categories.
[…]
Category of Question $Mean\pm SD$ $t$-value $p$-value 95% CI
ChatGPT CounterQuill
HS Identification Confidence 4.18$\pm$1.88 6.24$\pm$0.75 -4.280 ¡0.001*** [-3.079, -1.039]
Brainstorming Effectiveness 3.24$\pm$1.44 6.35$\pm$0.70 -8.849 ¡0.001*** [-3.865, -2.371]
Self-Efficacy in CS Writing 3.29$\pm$1.83 6.18$\pm$0.81 -7.356 ¡0.001*** [-3.713, -2.052]
Engagement with AI 3.59$\pm$1.73 6.00$\pm$0.87 -5.734 ¡0.001*** [-3.303, -1.520]
Satisfaction with CS 4.00$\pm$1.58 6.47$\pm$0.51 -7.668 ¡0.001*** [-3.154, -1.788]
Willingness to Post Online 3.53$\pm$1.46 6.29$\pm$0.77 -10.444 ¡0.001*** [-3.326, -2.204]
Figure 11. The violin plot of evaluating ChatGPT and CounterQuill impact on counterspeech writing based on our customized questions.
Our result found statistically significant differences (p ¡ 0.001) between CounterQuill and ChatGPT across various measures. Participants reported higher confidence in identifying online hate speech, found CounterQuill more effective for brainstorming and increasing self-efficacy in writing counterspeech, felt more engaged and motivated when using CounterQuill, were more satisfied with the counterspeech created using CounterQuill, and were more likely to post the co-written counterspeech online when using CounterQuill compared to ChatGPT.
Overall, the violin plot (Figure 11) visually represents the score distributions across the two conditions based on our six customized questions. Participants found CounterQuill more effective in enhancing their ability to identify hate speech, providing brainstorming assistance, increasing self-efficacy, fostering engagement, and leading to satisfaction with the final counterspeech. Moreover, participants were more likely to post the counterspeech co-created with CounterQuill on social media than ChatGPT.
6.2. Qualitative Findings of Learning with CounterQuill
6.2.1. Evolving Perceptions of Hate Speech: From Explicit to Implicit Forms
The learning session played a crucial role in shaping participants’ understanding of hate speech, particularly distinguishing between explicit and implicit forms. Many participants (19 — P1 - 13, P15 - P20) came into the study with assumptions about what does and does not constitute hate speech. While most (15 — P1 - P5, P7, P8, P10 - P13, P15 - P20) found it easy to recognize explicit hate speech, identifying implicit hate speech proved more challenging.
[…]
Participant 12: Now, I think implicit is harder to recognize because many people are not aware that what they’re saying is actually hate speech. They might think it’s just a normal thing to say, but there’s implicit[ly] hateful meaning behind it. So, implicit hate speech can be trickier to point out because it’s more subtle and social[ly] normalized.
6.2.2. Shifting Perceptions: Recognizing the Lasting Impact of Implicit Hate Speech
Prior to the learning session, participants had varying levels of awareness about the impact of hate speech. Some, like P19, initially believed that implicit hate speech was less harmful than explicit hate speech, stating, I think (implicit hate speech) is less harmful (than explicit hate speech), it’s not direct, probably doesn’t hurt too many people..
However, the learning session, which included explaining the impact of different types of hate speech and testing participants’ knowledge, helped to create a unified understanding and changed many participants’ perceptions. After completing the session, most participants (17 — P1 - P5, P7, P8, P10 - P13, P15 - P20) recognized that despite its subtlety, implicit hate speech could have a more lasting and harmful impact than explicit hate speech. This shift in perspective is exemplified by P5’s reflections following the learning session:
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=artifact, linewidth=0pt, backgroundcolor=artifact!10 {mdframed}
Participant 5: Before the learning session, I thought explicit hate speech was worse because it’s so obvious. But now I feel implicit [hate speech] has a much more lasting and harmful impact … The impact of implicit [hate speech] may be more rooted in society [be]cause only people [who belong to this] group can identify it as hate speech, [other people] may not think of it as hate speech.
Additionally, participants attributed the significant impact of implicit hate speech to several factors, including its pervasiveness (14 — P1, P4, P5, P7, P8, P10 - P13, P15-P19) and the difficulty of moderating it (8 — P2, P3, P10, P16 - P20).
6.2.3. Familiarizing with Counterspeech: From Awareness to Understanding
Our learning session introduced participants to the concept of counterspeech and its impact. Prior to the session, participants had varying levels of familiarity with counterspeech. Some participants (12 — P1 - P3, P8, P10 - P12, P14 - P17, P19) had a general understanding of counterspeech as a way to respond to hate speech. Two participants (P4, P13) had even posted responses to online hate speech prior to the session, but they were unaware that their behavior was considered counterspeech. Others (6 — P5, P6, P7, P9, P18, P20) were less familiar with the concept of counterspeech.
After completing the learning session, participants defined counterspeech as a respectful (18 — P2 - P12, P14 - P20) and non-escalating (8 — P1, P3, P6, P7, P9, P13, P18, P20) way to respond to hate speech. P4 provided a clear example of this understanding:
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=artifact, linewidth=0pt, backgroundcolor=artifact!10 {mdframed}
Participant 4: I did respond to some hate [speech] before, but I never knew the term (of that behavior) … Counterspeech is a respectful way to address hate speech. It addresses (the targets are) human, and we should treat each other kindly … It’s a way to respond to hate [without] more hate, but [with an attempt to] get some understanding.
6.2.4. Empathy-Based Counterspeech: Recognizing Its Effectiveness
Our learning session also introduced participants to the concept of empathy-based counterspeech. Prior to the session, few participants (4 — P2, P4, P6, P18) were aware of this specific approach to counterspeech. However, after completing the learning session, many participants (19 — P1 - P12, P14 - P20) recognized the effectiveness of empathy-based counterspeech in countering hate content.
[…]
Participant 6: I like what I learned empathy-based counterspeech (from the CounterQuill). If someone posted a hateful comment about gay people, I could respond like: “Hey, just so you know — my son is gay, and he’s a lot more than what you said in your comment. He has dreams, hopes, and feelings.” By sharing my story, I can encourage people to reconsider their views.
[…]
Participant 18: After the learning session, I had more to think about. … humorous counterspeech can still come across as implicitly hateful. For example, if someone posts a negative comment about body image and you respond with something like, “Well, you look ugly too,” it doesn’t really work. It’s just you’re joking back at them in a way of negativity. In my opinion, the best way to convey counterspeech is through empathy.
Overall, the learning session expanded participants’ understanding of counterspeech, particularly the value of empathy-based approaches in challenging hate speech effectively.
6.3. Qualitative Findings of Brainstorming with CounterQuill
6.3.1. Developing Hate Speech Identification Skills through Highlighting Practice
Our brainstorming session introduced participants to a color-coded highlighting practice, helping them identify the key components of hate speech: (1) Aspects of an individual or a group’s identity; (2) Action or perception that dehumanizes that person or group.
Prior to the highlighting practice, most participants (19 — P1 - 13, P15 - P20) had a general understanding of hate speech but struggled to articulate its specific components. As P11 mentioned, “I knew hate speech targeted certain groups and said bad things about them, but I couldn’t really break it down further than that.”.
However, after completing the highlighting practice, participants (17 — P1 - P13, P16, P17, P19, P20) found that the “the use of distinct colors (for targeted individuals or groups’ identities and dehumanizing actions) provided a clear visual breakdown of the hate speech components”, making it easier to grasp the core issues. P5 emphasized the benefits of this approach:
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=expert, linewidth=0pt, backgroundcolor=expert!10 {mdframed}
Participant 5: The highlighting made it easy to get the main points [in the hate speech]. When I read the example, I got what was being said and the negative things [that targeted the group]. It showed what the hate speech was about and the specific bad things [being said].
Through practicing highlighting, most participants (16 — P1 - P7, P9 - P13, P16, P17, P19, P20) gained a clearer understanding of how to identify targeted identities and dehumanizing language in real-world contexts. Some participants (8 — P4, P7, P8, P12, P13, P15, P19, P20) mentioned that it was their first time recognizing these key components of hate speech. P12 shared how the practice enhanced their attention to specific words and stereotypical behaviors:
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=expert, linewidth=0pt, backgroundcolor=expert!10 {mdframed}
Participant 12: After completing the highlighting task, I would start looking for words that try to point at particular groups based on their ethnicity or sexual orientation. And, I would pay more attention to small actions that people take could be stereotyped. I’ll look out for when identifying the key things (in hate speech).
In addition to the hands-on practice of highlighting the targeted identity and dehumanizing action, most participants (15 — P1 - P7, P9 - P12, P15, P17, P19, P20) also described benefiting from the system’s highlighting feedback and comparison features (§4.2). The system-generated suggestions helped participants correct and refine their understanding of targeted identities and dehumanizing language. P13 described how the feedback helped her identify problematic aspects they had initially overlooked:
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=expert, linewidth=0pt, backgroundcolor=expert!10 {mdframed}
Participant 13: After comparing my highlights with the system’s suggestions, … I see them as a good reference to better understand [the hatespeech]. It gave me a sense of additional aspects to consider, like the focus on women [as the target identity], not just independent women, and how even mentioning instinct is part of the hate speech. Those were elements I didn’t recognize.
For reference hate speech: “All of the independent women I knew eventually settled down and got married. Your instinct will kick in soon.”
Overall, the highlighting practice, coupled with the system’s feedback, effectively enhanced participants’ ability to identify and understand the key components of hate speech. As P6 summarized:
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=expert, linewidth=0pt, backgroundcolor=expert!10 {mdframed}
Participant 6: … Doing this highlight practice has really increased my confidence when it comes to finding targeted groups and [dehumanizing] language [in hate speech]. Before I tried this, I probably missed some of the more subtle stuff … Everyone knows hate speech is bad, but after highlighting it will let you understand why it’s so toxic.
6.3.2. Developing Empathetic Reflection through Reflective Questions
Before engaging in brainstorming process, participants’ initial reactions to hate speech varied. Some (8 — P3, P5, P8, P11, P13, P14, P18, P19) mentioned that they would typically respond with anger or frustration, while others (6 — P1, P4, P6, P7, P10, P17) expressed uncertainty about how to counter hate speech effectively. However, many participants (16 —P1, P2, P4, P6, P7 - P11, P13, P14 - P18, P20) noted that answering the questions in the Brainstorming session changed their approach to responding to hate speech compared to their initial reactions. For example, P8 commented:
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=expert, linewidth=0pt, backgroundcolor=expert!10 {mdframed}
Participant 8: When I was answering the first question, I was thinking about why this was wrong and trying to explain it. In the second part, I tried to show some empathy. So it changed how I approached it. If I didn’t answer those questions, I might have been more offensive in my response. But now I feel like I can write it more constructively, and say why it’s wrong and how the targeted people might feel.
Additionally, as mentioned by some participants (6 — P2, P7 - P9, P16, P20), the question-answering process guided them to engage in empathy-based counterspeech by considering the harmful impact and feelings of the targeted group. P2 elaborated on this shift in perspective:
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=expert, linewidth=0pt, backgroundcolor=expert!10 {mdframed}
Participant 2: … In the past when responding to or countering, I wasn’t asking myself questions like this internally, like what are the negative stereotypes, or considering the feelings of the targeted person and how it impacts them. It’s not something I intuitively think about. So I think this really leads me engage in that empathy-based counterspeech we talked about before. It brings my perspective on what’s happening, why it’s harmful, and what the impact is.
In addition to the questions-answering, most participants (16 — P1-P7, P9-P12, P15, P17, P19, P20) also described benefiting from the system’s feedback on their response reflection and suggested counterspeech strategies (§3.2 System Design - Brainstorming). These features helped reduce their anxiety and mental load, particularly when they were unsure how to begin writing counterspeech. P5 specifically noted that the system’s suggestions alleviated the stress of starting counterspeech from scratch.
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=expert, linewidth=0pt, backgroundcolor=expert!10 {mdframed}
Participant 5: After I got the feedback, there were things I wanted to add to my comments or input. It was like, ’Oh, I could add that.’ The feedback gave me a good starting point for writing counterspeech. If you asked me to write a counterspeech from scratch, I would start from a place of uncertainty, and it might not be very straightforward. But the feedback gave me a clear target to move on.
Some participants (12 — P1-P6, P9-P12, P19, P20) noted that the system’s feedback was essential in helping them think in-depth about the key elements of the hate speech they were analyzing, which guided their approach to crafting a response. P20, for example, detailed:
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=expert, linewidth=0pt, backgroundcolor=expert!10 {mdframed}
Participant 20: I think the feedback is good because, firstly, the feedback seems to respond to things that I said. So, they identify points that are useful for writing counterspeech. But then later, it also expands more on [either] why what I said is relevant [or] how it can be used in counterspeech, and how that can lead to different kinds of responses to counterspeech. So, I think the format and the personalized response [by the system] are helpful. Yeah, it’s very detailed and valuable.
6.4. Qualitative Findings of Writing with CounterQuill
6.4.1. Empowering Users through CounterQuill’s Writing Session: From Brainstorming to Confident Counterspeech
Most participants commented on the intuitive design and user-friendliness of CounterQuill’s writing interface. The friendly layout and convenient features allowed participants to focus on composing their counterspeech without being hindered by complex technology. P2 summarized his experience:
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=training, linewidth=0pt, backgroundcolor=training!10 {mdframed}
Participant 2: The writing interface is very user-friendly and easy to use. It allowed me to concentrate on writing my response without getting distracted by complicated features.
[…]
Participant 5: Before using CounterQuill, I wasn’t sure where to start when writing counterspeech. I had ideas, but I didn’t know how to structure them or make them sound convincing.
However, after using CounterQuill, participants reported that the integration of brainstorming notes into the writing interface provided a smooth transition from the ideation stage to the actual writing process. Having the ability to refer to these notes during writing was beneficial (11 — P2, P5, P6, P10, P11, P12, P14, P17 - P20), as it helped participants structure their responses and include key points. P18 elaborated:
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=training, linewidth=0pt, backgroundcolor=training!10 {mdframed}
Participant 18: … looking back at the notes was definitely helpful. I tried to structure my entire response based on the notes and the main points I had taken in the earlier sessions. Breaking things down in the notes beforehand really did help me in forming the text for my counterspeech … It just made the process a lot smoother.
Moreover, using the participants’ responses from the brainstorming session as an initial draft for the counterspeech reduced their psychological pressure and cognitive load of starting from scratch (14 — P2 - P6, P9, P11, P12, P14, P15, P17 - P20). This approach also improved participants’ confidence and helped them effectively capture the key points they wanted to address (5 — P2, P8, P10, P11, P12). P8 noted how this feature significantly increased their confidence.
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=training, linewidth=0pt, backgroundcolor=training!10 {mdframed}
Participant 8: … Starting a draft based on my own ideas or [responses] made the writing [process] less stressful. I felt like I had an initial point, and it increased my confidence [for writing counterspeech] and now I know where I need to start about the counterspeech.
6.4.2. Co-writing with CounterQuill: Selective Rewriting, Improved Confidence, and Increased Ownership
Once participants completed the draft of their counterspeech, they could select specific content and request writing assistance from CounterQuill. Most participants (16 — P1, P3, P5 - P11, P14, P15 - P20) found that the fine-grained control over choosing particular sentences or paragraphs for AI-assisted polishing was more personalized than current conversational AI (i.e., ChatGPT), providing a level of customization that other tools don’t offer. For example, P1 emphasized this advantage:
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=training, linewidth=0pt, backgroundcolor=training!10 {mdframed}
Participant 1: I think the ability to select specific sentences for rewriting is valuable. I wrote a pretty good draft and I want to keep most of it, but maybe there are one or two sentences that I want to include and I don’t really know how to phrase them. Having this system in place is helpful because you can keep what you wrote originally while also editing it part by part with the AI, all within the context of the full response.
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=training, linewidth=0pt, backgroundcolor=training!10 {mdframed}
Participant 1: … I think that’s a lot better than ChatGPT because if you want to do selective edits with ChatGPT, you have to tell it specifically that you don’t want to change certain parts, or you would have to take those parts out and then only insert the part that you want to edit. But with this tool, you can just do the selective rewriting in one place, which is an improvement.
In addition to selective rewriting, participants mentioned the features of CounterQuill’s writing assistant. Among these, many participants (12 — P2 - P6, P10, P11, P12, P17 - P20) reported that grammar checking was the first feature they would use. All participants (20) had thoughts on the empathetic rewriting feature: the empathetic rephrasing suggestions helped participants better express their ideas (17 — P1 - P12, P15 - P19), making their responses more relatable and humanizing (17 — P1 - P8, P10 - P13, P15 - P18, P20). The empathetic rephrasing suggestions helped participants better articulate their thoughts (11 — P2, P6 - P13, P19, P20), enhanced the impact of their counterspeech (15 — P1, P2, P3 - P9, P11 - P14, P18, P20) and give them emotional support (7 — P2, P3, P5, P7, P12, P14, P18). For example, P2 felt more connected and less confrontational in writing counterspeech and shared how the system improved the original counterspeech:
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=training, linewidth=0pt, backgroundcolor=training!10 {mdframed}
Participant 2: I think using this system to polish and refine my original draft really helped because it made my comment a little shorter but also made it more empathetic. It generally kept my content the same. The AI also added their own touch at the very end, something like ‘This is an opportunity for us to work together.’ I think that really added a sense of empathy and encouragement for mutual understanding. It doesn’t seem like you’re talking at them but talking with them.”
[…]
Participant 7: CounterQuill’s selective rewrite is helpful because I can keep what I wrote originally while also editing it part by part together with the AI … and the final response is created by CounterQuill and me together.
[…]
Participant 14: I tried many prompts (with ChatGPT), but even if it gave me 10 different counterspeech, I wouldn’t use any of them … because they are too AI-like.
In contrast, some participants expressed a preference for continuing to use CounterQuill’s writing features in future real-world online interactions. They particularly appreciated how CounterQuill integrated prompts like tone adjustment and rewriting into convenient buttons, which made the counterspeech writing process more efficient than using other chatbots (“I am likely to use CounterQuill in the future because its system features, like changing tone, help me write counterspeech more quickly than some chatbots” - P2).
7. DISCUSSION
7.1. Design Implications for AI-Mediated Counterspeech Tools
Designing AI tools that cater to a wide range of individuals is essential, particularly in social media environments where diverse attitudes towards counterspeech are prevalent. Prior work has highlighted the varying motivations, barriers, and skills among different groups when countering online hate speech (Mun et al., 2024; Ping et al., 2024; Zhu and Bhat, 2021). This diversity prompts the question: how can we design customized AI tools that effectively function in social media environments, where users with diverse attitudes engage in counterspeech? How can these tools influence the thought patterns and behaviors of such varied groups?
Our study provides insights to help address these questions in the context of improving ways to engage in effective counterspeech. Our findings revealed varied attitudes from participants when faced with the challenge of writing counterspeech. For instance, P1 expressed a strong moral obligation to write counterspeech (‘well I think about it this way, if someone is saying something nasty and you are silent, it’s like letting the bad stuff win. I just feel like engaging in counterspeech is the right thing to do.’ ), while P4 hesitated to participate due to fear of potential backlash or lack of confidence in writing abilities. For instance, when asked about how confident they are in writing counterspeech before using CounterQuill, P4 expressed concerns— ‘I am always worried about what might happen if I speak back. What if people attack me for it? Also, I’m not sure if I can speak my thoughts well enough.’ Such varying attitudes and reactions reinforce us to design AI-assisted tools that can provide tailored suggestions to users, aligning with their unique comfort levels and engagement preferences in countering hate speech. Future research could explore the potential of embedding tools like CounterQuill directly within social media platforms (e.g., Facebook, Instagram, Reddit, Twitter) as extensions for browsers. Such an approach could make counterspeech tools more accessible and convenient for users, reducing the time spent navigating web interfaces and allowing for real-time, context-sensitive suggestions. By focusing on the balance between time and outcomes, integrating features like tone adjustment and rewriting prompts may provide users with immediate, relevant guidance based on their interaction history, such as posts and comments.
Moreover, as P5 stated - “The highlighting exercise made it easy to get the main points [in the hate speech]”, users can benefit from AI offering more personalized explanations. For instance, as our findings show, these explanations can focus on specific elements like the identity of the targeted individual/ group or the dehumanizing actions to help users comprehend and respond to hate speech effectively.
In addition, participants were inconsistent in their understanding of counterspeech. For example, participants (P4, P18) recognized their limitations in responding to hate speech but were not aware of specific terminology (‘I did respond to some hate [speech] before, but I never [knew] the term (counterspeech)’), while others were unaware of the terminology altogether (P5, P7). This disparity in user perceptions extended to their goals and motivations in engaging in counterspeech. Therefore, AI-assisted counterspeech research should address users’ varied needs and backgrounds engaging with these tools, particularly LLM-powered systems. To enhance counterspeech efforts, these tools should prioritize adaptability and personalization, striking a balance between efficiency and effectiveness. AI-mediated tools should aim to provide tailored guidance based on users’ personal experiences and specific goals in engaging with counterspeech.
7.2. AI-Mediated Communication for Counterspeech Ownership & Engagement
With the advancement of AI technology, AI-mediated communication (AIMC) offers new possibilities for assisting individuals in overcoming these barriers and writing counterspeech (Mun et al., 2024; Zhu and Bhat, 2021; Ping et al., 2024). However, the introduction of AIMC has also raised concerns, particularly regarding participants’ sense of ownership of the AI-generated content, which is related to their engagement with counterspeech (e.g., whether or not to post counterspeech on social media).
In this section, we discuss potential outcomes that may arise from AI-mediated counterspeech, such as participants’ sense of ownership over the co-written counterspeech and their level of engagement.
Figure 12. The relationship between user effort, sense of ownership, and willingness to post counterspeech on social media: more cognitive effort invested in the AI-mediated system leads to a greater sense of ownership and increased likelihood of posting the co-created content online. The blue background area is indicated: past research has demonstrated the correlation between cognitive effort and ownership of human AI co-created content.
7.2.1. How The Sense of Ownership Manifested through Counterquill Usage
Our findings discovered that when participants engage in collaborative writing with CounterQuill, they develop a sense of ownership over the counterspeech. This occurs when participants incorporate personal experiences and values into their writing process and make revisions based on CounterQuill’s suggestions. Our findings echo previous HCI research, (Figure 12), showing that the more meaningful cognitive effort people invest in collaborating with AIMC systems, the stronger their sense of ownership becomes (Louie et al., 2021, 2020a; Pierce et al., 2003; Mieczkowski, 2022).
Through our collaboration with CounterQuill, we found that participants were surprised by how the system co-authored counterspeech with them. Many referred to CounterQuill as a writing “partner” or “friend” rather than simply an AI or model. We believe the following features of the system contributed to the increased sense of ownership. First, participants could maintain control by selecting AI-generated text suggestions or querying the system for tone modifications using the “AI Writing Assistant” interface. Second, CounterQuill preserved their efforts, unique voices, and tones. The system’s generated results maintained the context and semantics of the user’s original content, aligning with their intentions and personal experiences. Last, participants in the human-AI collaboration condition could actively insert and delete personal content during the writing process. They engaged in thoughtful deliberation while writing and reported the feeling of initiative and control over the writing process, leading to a stronger sense of ownership.
In addition, we found that the collaborative nature of the co-writing process promoted a sense of shared goals, echoing Caspi et al.’s findings. (Caspi and Blau, 2011). Participants perceived the final counterspeech as “ours” rather than merely “mine,” further enhancing their sense of ownership. Specifically, many participants in our study believed that the final response was completed by “CounterQuill and me together.” This co-ownership emerged when participants invested cognitive effort, energy, and even personal values, and the AI system provided suggestions and guidance that aligned with the participants’ experiences and personal values.
7.2.2. Comparing ChatGPT and CounterQuill : Ownership and Engagement
However, the sense of ownership over information is not without complexity. People tend to believe that the results of co-creation with AI belong to “us”, while simultaneously acknowledging that AI-generated text is trained from the language of potentially billions of others. This tension, as suggested by Mieczkowski (2022), may lead to a slight defensiveness in some individuals due to the perceived threat to their sense of ownership (Mieczkowski, 2022).
Through our study, we discovered that participants’ sense of ownership over counterspeech was diminished during the baseline task. Most participants pointed out that “CounterQuill is different from ChatGPT,” because people typically invest less time and interaction with conversational AI tools (i.e., ChatGPT), resulting in a lower sense of ownership. They often describe these tools as “assistants” rather than “collaborators”. This finding echoes Louie et al. and Mieczkowski’s study (Figure 12), where people expressed a lower sense of ownership due to their limited investment in the task (Mieczkowski, 2022; Louie et al., 2020a), even describing themselves as “lazy” (Louie et al., 2020a, b).
Furthermore, our study found that during interactions with ChatGPT, participants felt “passive” (Dummer et al., 2009; Synofzik et al., 2008) because they could neither predict the output nor change it after insertion. Consequently, participants opted for “lazy prompts” when conversing with ChatGPT. However, even when individual participants attempted to write elaborate prompts, they experienced resistance (“I tried many prompts, but even if it gave me 10 different counterspeech, I wouldn’t use any of them, because they are too AI-like” - P14). These findings are consistent with previous studies (Louie et al., 2020a; Kim and Song, 2023), indicating that human interactions with conversational AI, such as ChatGPT, lack the back-and-forth exchange and the investment of personal effort. Consequently, people generally express a weak sense of ownership over the results generated by these AI “assistants” and refrain from using them.
7.2.3. The Relationship Between Ownership and Willingness to Share
Our study reveals that participants’ perception of ownership and the amount of time (perceived effort) they invest in CounterQuill influence their end goal of whether they will post the counterspeech online.
Our quantitative analysis (§6.1) showed that collaborating with ChatGPT required less cognitive effort than CounterQuill. However, in our customized questions, participants reported higher satisfaction with the final counterspeech generated in collaboration with CounterQuill, with an average user rating of 6.47, significantly higher than ChatGPT (4.00) by 62%. Moreover, participants were more inclined to publish the final generated content online when collaborating with CounterQuill, with an average user rating of 6.29, considerably higher than ChatGPT (3.53) by 78%. Although our user study did not directly measure perceived ownership of the generated content, we found that participants’ satisfaction and willingness to publish online, which can be seen as a proxy of perceived ownership, were higher in CounterQuill’s condition. This finding echoes previous research by Lee et al. (2015), who discovered that people are willing to share publicly for various motivations, for example, sharing their own ideas and speeches (“To express my actual self” and “To share my personal information with others”) (Lee et al., 2015).
Combining our qualitative analysis, our findings indicate that when working with AI, people tend to develop a strong sense of ownership over co-created goals due to the investment of time, energy, and even personal values required in the process of achieving these goals. Consequently, they are more willing to publish content that belongs to ”ours” on social media. However, when AI-generated content is produced quickly and with minimal effort, the sense of ownership loss becomes particularly evident. Furthermore, when users consider the potential perspectives of others - how the recipients of the information might view the sender’s use of AI-generated text - they are less likely to publish it on social media.
8. CONCLUSION
In this work, we introduced CounterQuill, an AI-mediated system designed to help users write effective counterspeech. CounterQuill guides users through a learning session to understand hate speech and counterspeech concepts, a brainstorming session to identify hateful elements and generate response ideas, and a co-writing session to draft and refine their counterspeech.
Our user study revealed that CounterQuill’s learning session effectively educated users about hate speech and empathy-based counterspeech strategies. The brainstorming session, with its AI guidance and feedback, improved users’ confidence and perceived ability to engage in counterspeech. Furthermore, the co-writing session, which provided AI writing assistance and personalization options, enhanced users’ self-efficacy and ownership of the counterspeech they created. Interestingly, we found that the more meaningful effort users invested in collaborating with CounterQuill, the stronger their sense of ownership over the co-created content, leading to a higher likelihood of posting the counterspeech on social media.
Limitations and future work. While our research provides insights into the potential of using AI-mediated system for countering online hate speech, further research is needed to explore these nuances of human-AI collaboration and the factors that influence users’ willingness to participate. For example, past research found that individuals who confronted those who were engaging in hate speech often became targets themselves (Costello et al., 2017; Hawdon et al., 2017). This victimization, in turn, can reduce one’s willingness to participate if the person sees hate speech in the future (Hawdon and Costello, 2022). One limitation of our user study is that participants’ prior experience and competence with AI and counterspeech may have influenced their performance and perceptions when using CounterQuill and ChatGPT, potentially impacting the study’s results. Additionally, our study was restricted to participants responding to a single hate post. Future research should include multiple hate posts to gain more comprehensive insights and better assess CounterQuill’s effectiveness across different contexts.
Future research should investigate if using AI-generated counterspeech reduces the likelihood of the user being targeted or provides the user greater confidence to generate counterspeech even if they are victimized. If human-AI collaboration can both improve users’ sense of ownership and reduce their apprehension about participating in counterspeech, we will be able to design more effective tools to combat online hate speech and foster healthier online communities.
Building on this, we propose developing CounterQuill as a Chrome or Edge extension that can be integrated directly into social media platforms like Reddit and Twitter. Our future work will include a long-term user study of the CounterQuill extension, where participants will use it in their daily lives to identify hate speech and collaborate with AI to compose counterspeech responses on real social media platforms. We also plan to conduct more granular comparisons to showcase the unique writing features of the CounterQuill extension with other AI tools.
Appendix A pre-screening survey
  1. (1)
    Have you used Large Language Model (LLM) tools or an AI chatbot, such as ChatGPT, Google Gemini or Claude?
    1. (a)
      Yes
    2. (b)
      No
  2. (2)
    You have indicated that you have used an AI chatbot, such as ChatGPT. Can you briefly describe this experience?
  3. (3)
    Have you ever countered hate content on online social media?
    1. (a)
      Yes
    2. (b)
      No
  4. (4)
    You have indicated that you have countered hate speech on social media. Can you briefly describe this experience?
  5. (5)
    What is your age?
  6. (6)
    What is your gender identity? Select as many that apply.
    1. (a)
      Woman
    2. (b)
      Man
    3. (c)
      Trans man
    4. (d)
      Trans woman
    5. (e)
      Genderqueer
    6. (f)
      Agender
    7. (g)
      Autigender
    8. (h)
      Prefer not to answer
    9. (i)
      Other
  7. (7)
    What is your ethnicity?
    1. (a)
      Asian/Asian American
    2. (b)
      Black/African American
    3. (c)
      Hispanic/Latina/o/x
    4. (d)
      Middle Eastern/North African
    5. (e)
      Native American/American Indian
    6. (f)
      Native Hawaiian/Pacific Islander
    7. (g)
      White/European American
    8. (h)
      Other
  8. (8)
    What is your education level?
    1. (a)
      Less than high school
    2. (b)
      High school graduate
    3. (c)
      Some college
    4. (d)
      Associate degree
    5. (e)
      Bachelor’s degree
    6. (f)
      Some postgraduate work
    7. (g)
      Master’s degree
    8. (h)
      Doctorate degree
  9. (9)
    What is your sexual orientation?
    1. (a)
      Heterosexual
    2. (b)
      Homosexual
    3. (c)
      Bisexual
    4. (d)
      Prefer not to say
    5. (e)
      Other
  10. (10)
    In general, how would you describe your political views?
    1. (a)
      Very conservative
    2. (b)
      Conservative
    3. (c)
      Moderate
    4. (d)
      Liberal
    5. (e)
      Very Liberal
  11. (11)
    We would like to invite you to a follow-up interview study based on your responses. Are you interested in participating?
    1. (a)
      Yes
    2. (b)
      No
  12. (12)
    Are you comfortable speaking and participating in this study in English?
    1. (a)
      Yes
    2. (b)
      No
  13. (13)
    Do you have a laptop/PC?
    1. (a)
      Yes
    2. (b)
      No
Appendix B Guiding interview questions
  1. (1)
    How do you perceive implicit hate speech after reviewing the learning materials provided?
  2. (2)
    What are your thoughts on the negative impacts of explicit versus implicit hate speech?
  3. (3)
    How would you describe empathy-based counterspeech, and what role do you think it plays in addressing hate speech?
  4. (4)
    What were your impressions of using colors to highlight targeted identities and dehumanizing actions? How did this influence your understanding?
  5. (5)
    Do you feel that your ability to identify targeted identities and dehumanizing language has evolved through this exercise? If so, in what ways?
  6. (6)
    How did the brainstorming questions influence your approach to responding to hate speech compared to your initial reaction?
  7. (7)
    How did you find the personalized feedback provided by the system? In what ways did it help or hinder your process?
  8. (8)
    How did you perceive the AI’s ”polish” and ”rewrite as empathetic” features in refining your response?
  9. (9)
    How do you value the ability to selectively rewrite specific sentences using AI compared to other editing tools?
  10. (10)
    How did the combination of learning, brainstorming, and writing aids affect your confidence in crafting counterspeech?
Appendix C post-study questionnaire
To save space on the page, we have combined the post-study questions, as the content of the questions is identical, with the only difference being that Condition A refers to GPT and Condition B refers to CounterQuill.
Please rate your experience using ChatGPT (condition A) / CounterQuill (condition B) to write counter-speech on the following dimensions. For each dimension, select a number from 1 (very low) to 7 (very high).
  1. (1)
    Please enter your participation number.
  2. (2)
    How much mental activity was required to use ChatGPT (condition A) / CounterQuill (condition B) for counter-speech writing (e.g., how much thinking or deciding was required to perform the task)?
    Very Low
     1  2  3  4  5  6  7  Very High
  3. (3)
    How much physical activity was required to use ChatGPT (condition A) / CounterQuill (condition B) for counter-speech writing (e.g., typing, clicking, etc.)?
    Very Low
     1  2  3  4  5  6  7  Very High
  4. (4)
    How much time pressure did you feel due to the pace at which the task of counter-speech writing using ChatGPT (condition A) / CounterQuill (condition B) occurred?
    Very Low
     1  2  3  4  5  6  7  Very High
  5. (5)
    How hard did you have to work (mentally and physically) to accomplish your level of performance in using ChatGPT (condition A) / CounterQuill (condition B) for counter-speech writing?
    Very Low
     1  2  3  4  5  6  7  Very High
  6. (6)
    How successful do you think you were in accomplishing the goals of the task (writing effective counter-speech using ChatGPT (condition A) / CounterQuill (condition B))?
    Very Low
     1  2  3  4  5  6  7  Very High
  7. (7)
    How insecure, discouraged, irritated, stressed, and annoyed versus secure, gratified, content, relaxed, and complacent did you feel during the task of using ChatGPT (condition A) / CounterQuill (condition B) for counter-speech writing?
    Very Low
     1  2  3  4  5  6  7  Very High
Please rate your experience using ChatGPT (condition A) / CounterQuill (condition B) to write counter-speech on the following dimensions. For each question, select a number from 1 (strongly disagree) to 7 (strongly agree).
  1. (1)
    How confident are you in your ability to identify online hate speech that dehumanizes individuals or groups based on their identity after co-writing with ChatGPT (condition A) / CounterQuill (condition B)?
    Strongly Disagree
     1  2  3  4  5  6  7  Strongly Agree
  2. (2)
    How effective is ChatGPT (condition A) / CounterQuill (condition B) at helping you with step-by-step guides for planning and thinking about counter-speech writing?
    Strongly Disagree
     1  2  3  4  5  6  7  Strongly Agree
  3. (3)
    How confident do you feel about the quality of the counter-speech you wrote with the help of ChatGPT (condition A) / CounterQuill (condition B)?
    Strongly Disagree
     1  2  3  4  5  6  7  Strongly Agree
  4. (4)
    To what extent did ChatGPT (condition A) / CounterQuill (condition B) provide customized assistance that was helpful for your specific counter-speech writing needs?
    Not at all
     1  2  3  4  5  6  7  To a great extent
  5. (5)
    To what extent did using ChatGPT (condition A) / CounterQuill (condition B) increase your self-efficacy (belief in your own abilities) in writing effective counter-speech?
    Not at all
     1  2  3  4  5  6  7  To a great extent
  6. (6)
    How engaged and motivated did you feel while using ChatGPT (condition A) / CounterQuill (condition B) to write counter-speech?
    Not at all engaged and motivated
     1  2  3  4  5  6  7  Extremely engaged and motivated
  7. (7)
    How satisfied are you with the final counter-speech you created using ChatGPT (condition A) / CounterQuill (condition B)?
    Not at all satisfied
     1  2  3  4  5  6  7  Extremely satisfied
  8. (8)
    If you had the opportunity, how likely would you be to post this counter-speech created using ChatGPT (condition A) / CounterQuill (condition B) on social media?
    Not at all likely
     1  2  3  4  5  6  7  Extremely likely
Appendix D quiz questions
  • Q1.
    Which statement accurately reflects the nature of hate speech?
    • A.
      Hate speech primarily targets a person’s actions or choices.
    • B.
      Hate speech is any form of speech that someone finds offensive.
    • C.
      Hate speech specifically targets core aspects of a person’s or group’s inherent identity, such as race or religion.
    • D.
      Hate speech is the same as constructive criticism but with a negative tone.
  • Q2.
    Which of the following is NOT a consequence of implicit hate speech?
    • A.
      Desensitization to derogatory language, resulting in reduced empathy for victims and insensitivity towards hate speech.
    • B.
      Promoting important and constructive diverse viewpoints, fostering understanding and tolerance among different groups.
    • C.
      Inducing psychological and emotional harm, increasing feelings of fear, anger, and helplessness for individuals targeted by implicit hate speech.
    • D.
      Promoting prejudice and fostering bigotry and intolerance through stereotyping and negative perceptions of certain groups.
  • Q3.
    Which statement does not accurately reflect the role of counterspeech?
    • A.
      Counterspeech can be risky (e.g., leading to backlash).
    • B.
      Counterspeech can help people to change their hateful perspectives. .
    • C.
      Counterspeech can encourage others to also counter hate.
    • D.
      Counterspeech can always exacerbate tensions and escalate the conflict.
  • Q4.
    Which of the following statements about empathy-based counterspeech aligns with the given learning materials?
    • A.
      Empathy-based counterspeech is most effective when it involves retaliatory attacks against the attacker.
    • B.
      Empathy-based counterspeech humanizes the victim and highlights the potential harm caused by the attacker’s behavior.
    • C.
      Empathy-based counterspeech is ineffective unless it includes aggressive language and strong condemnations.
    • D.
      Empathy-based counterspeech is only effective when delivered by a single user rather than a collective group.
Appendix E CounterQuill prompt
Because CounterQuill prompts are long, we created a GitHub repository to save our prompts. See https://github.com/AnonymousUseOnly/Group25 or Supplementary Materials for details.
Appendix F Examples of Hate Speech
Warning: Table A1 lists examples of hate speech used by user study in order of participants. The table contains speech examples that may be offensive or upsetting. These examples are presented solely for academic and research purposes and do not reflect the views of the author.
Table A1. Examples of Hate Speech and Associated Topics
Topic Example
#gender Women are naturally better at caring for others and doing housework than men, who are a mess and illogical at work.
#disability ”Mentally disabled people should not be allowed to vote. They are a burden to society.”
#race ”I saw a random black man jogging nearby our house today and now I feel unsafe walking around my own neighborhood.”
#sexual_orient Lesbians are often stereotyped as ‘tomboys,’ rejecting traditional female roles and appearances, and when I see them, I get uncomfortable because they’re not feminine at all.
#race I guessed you must be Japanese because in my experience Japanese people have very modern haircuts and clothing. Chinese people usually have very traditional looking clothing and hair, and you don’t look that at all! You look so chic and modern!
#gender ”It’s impressive how you handle such a demanding job and still manage to keep your family together. I guess it’s rare to find women who can balance both without letting one slip.”
#race ”Indeed, you are doing fine in a country with a majority white population. How is Mexico faring these days? People come here because Mexico couldn’t develop properly.”
#race It’s amazing how those kids from the inner city manage to graduate high school. I mean, considering their circumstances, you’d expect them not to achieve much.
#disability ”Persons with learning difficulties should be discouraged to attend primary school and they should be excluded from education.”
#sexual_orient ”Stop tucking your hair into your hat, you’ll attract a lot of lesbians that will want to be your girlfriend.”
#gender ”Women are often seen as being too emotional to handle leadership roles effectively.”
#race ”I’ve noticed that neighborhoods tend to go downhill when more of black people move in, you know? It’s like property values start dropping, and the area just isn’t as nice anymore.”
#disability ”Children with Down Syndrome should be isolated, because in any case normal children would make them feel different, ending up in unwanted discrimination.”
#sexual_orient ”Gays are sick. They should be isolated and isolated soon. They are a danger to society.”
#religion ”Islamists/Muslims have mastered the corrupt art of playing the victim and blaming other cultures for the violent beliefs taught in their main books.”
#religion ”Right. I’m questioning if this was a legitimate event, though. It looks like a hoax perpetrated by the media, which is perceived to be owned predominantly by Jewish interests, to be honest.”
#religion ”Every time I board a plane and see someone wearing a turban, I can’t help but feel a bit nervous, as if the safety of the flight is suddenly in question. It’s just an instinctive reaction, I guess.”
#race ”It was cleaner when white people lived here. Then the Blacks came in and now look! There’s trash all over the streets, dogs running loose, and the smell of smoke every sunny day.”
Appendix G Results of Welch’s Two-Sample t-tests
We conducted Welch’s Two-Sample t-tests that accounts for unequal variances between groups (Sakai, 2016), as detailed in Tables A2 and A3. The results confirmed that the sequence—whether participants began with ChatGPT or CounterQuill—did not significantly influence their performance or feedback on the system.
Table A2. Results of Welch’s Two-Sample t-tests assessing the effect of counterbalancing the order of system usage (ChatGPT First vs. CounterQuill First) on NASA-TLX scores. The analysis was conducted to determine if the sequence in which participants used ChatGPT and CounterQuill influenced their perceived workload across various dimensions (Mental Demand, Physical Demand, Temporal Demand, Performance, Effort, and Frustration). No significant differences were found (all p ¿ 0.05), indicating that the order of system usage did not affect participants’ performance or feedback.
Platform NASA-TLX Items $Mean\pm SD$ t-value p-value 95% CI
ChatGPT First CounterQuill First
ChatGPT Mental Demand 4.80$\pm$1.32 4.30$\pm$1.95 0.673 0.511 [-1.077, 2.077]
Physical Demand 2.40$\pm$1.65 1.90$\pm$1.52 0.705 0.490 [-0.991, 1.991]
Temporal Demand 2.90$\pm$1.37 2.30$\pm$1.49 0.936 0.362 [-0.748, 1.948]
Performance 4.00$\pm$1.05 3.80$\pm$1.14 0.408 0.688 [-0.830, 1.230]
Effort 2.60$\pm$1.17 2.20$\pm$0.92 0.849 0.408 [-0.594, 1.394]
Frustration 3.40$\pm$1.90 2.60$\pm$2.17 0.878 0.392 [-1.118, 2.718]
CounterQuill Mental Demand 5.50$\pm$0.53 4.50$\pm$1.65 1.826 0.096 [-0.208, 2.208]
Physical Demand 3.90$\pm$1.66 2.80$\pm$1.69 1.468 0.159 [-0.474, 2.674]
Temporal Demand 4.90$\pm$1.73 4.10$\pm$1.20 1.203 0.246 [-0.610, 2.210]
Performance 6.20$\pm$0.79 6.50$\pm$0.71 -0.896 0.382 [-1.004, 0.404]
Effort 4.20$\pm$1.55 3.90$\pm$1.79 0.400 0.694 [-1.276, 1.876]
Frustration 1.70$\pm$0.82 1.30$\pm$0.67 1.188 0.251 [-0.309, 1.109]
Table A3. Results of Welch’s Two-Sample t-tests assessing the impact of counterbalancing the order of system usage (ChatGPT First vs. CounterQuill First) on responses to customized questions. The questions evaluated various aspects including Hate Speech (HS) Identification Confidence, Brainstorming Effectiveness, Self-Efficacy in Counterspeech (CS) Writing, Engagement with AI, Satisfaction with Counterspeech, and Willingness to Post Online. The analysis showed no significant differences in responses based on the order of system usage (all p ¿ 0.05), suggesting that the sequence in which participants used ChatGPT and CounterQuill did not influence their responses to these customized questions.
Platform Category of Question $Mean\pm SD$ $t$-value $p$-value 95% CI
ChatGPT First CounterQuill First
ChatGPT HS Identification Confidence 5.00$\pm$1.41 3.40$\pm$1.96 2.097 0.052 [-0.014, 3.214]
Brainstorming Effectiveness 3.50$\pm$1.27 3.00$\pm$1.49 0.808 0.430 [-0.803, 1.803]
Self-Efficacy in CS Writing 4.00$\pm$1.76 2.60$\pm$1.65 1.835 0.083 [-0.204, 3.004]
Engagement with AI 4.10$\pm$1.52 3.00$\pm$1.70 1.524 0.145 [-0.418, 2.618]
Satisfaction with CS 4.50$\pm$1.43 3.30$\pm$1.42 1.882 0.076 [-0.140, 2.540]
Willingness to Post Online 4.10$\pm$1.60 2.80$\pm$1.14 2.100 0.052 [-0.011, 2.611]
CounterQuill HS Identification Confidence 6.40$\pm$0.70 6.20$\pm$0.79 0.600 0.556 [-0.501, 0.901]
Brainstorming Effectiveness 6.60$\pm$0.52 6.20$\pm$0.79 1.342 0.199 [-0.234, 1.034]
Self-Efficacy in CS Writing 6.40$\pm$0.97 6.10$\pm$0.57 0.847 0.411 [-0.457, 1.057]
Engagement with AI 6.00$\pm$0.94 5.90$\pm$0.74 0.264 0.795 [-0.699, 0.899]
Satisfaction with CS 6.60$\pm$0.52 6.30$\pm$0.48 1.342 0.196 [-0.170, 0.770]
Willingness to Post Online 6.30$\pm$0.82 6.20$\pm$0.79 0.277 0.785 [-0.658, 0.858]