### v2

# CounterQuill: Investigating the Potential of Human-AI Collaboration in Online Counterspeech Writing

Abstract.
Online hate speech has become increasingly prevalent on social media, causing harm to individuals and society. While content moderation has received considerable attention, user-driven counterspeech remains a less explored yet promising approach. However, many people face fear of retaliation and difficulties in articulating effective responses. We introduce CounterQuill, an AI-mediated system that supports users in composing empathetic counterspeech. CounterQuill’s three-step workflow includes: (1) a learning session on hate speech and counterspeech; (2) a brainstorming session to identify key elements of hateful content and develop strategies; and (3) a co-writing session to refine counterspeech. In a within-subjects user study (N=20) comparing CounterQuill to ChatGPT, participants found CounterQuill’s structured guidance and collaboration more empowering, increasing their willingness to share their co-written counterspeech online. We discuss implications for designing AI-assisted, user-centered tools to mitigate online hate while preserving individual voices and a sense of ownership.
Human-AI Collaboration, counterspeech, large language model, writing assistants.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
Figure 1. CounterQuill’s three-stage workflow: (A) Learning Session builds users’ understanding of hate speech, counterspeech, and their broader implications through interactive lessons and quizzes; (B) Brainstorming Session guides users in identifying targeted identities and harmful assumptions, then brainstorming counterspeech strategies; and (C) Co-Writing Sessionoffers an AI-assisted environment for drafting, refining, and personalizing counterspeech with grammar checks, empathetic tone suggestions, and user-driven input.
1. Introduction
Starting last year and intensifying in the last few weeks under the new U.S. administration, many social media companies have begun lifting restrictions on hate speech (Dodds, 2025). Rather than relying on top-down platform moderation, sites like Facebook are increasingly shifting to a community-driven content moderation strategy, placing users in charge of identifying and addressing potentially false or offensive content—including hate speech (Meta, 2025). This shift coincides with a broader rise in harmful online content: the rapid growth of communication platforms has led to an overall increase in hate speech, causing significant harm to individuals and society (Holt and Bossler, 2020). A study by the Pew Research Center found that 41% of U.S. adults have experienced online hate speech, with 18% subjected to severe forms such as physical threats, persistent harassment, or sexual harassment (DUGGAN, 2017). Youth and young adults are even more vulnerable, with a cross-national study reporting that approximately 70% of individuals aged 18–25 have encountered hateful materials (Reichelmann et al., 2021).
Counterspeech is broadly defined as communication intended to counteract potential harm caused by other speech (Cepollaro et al., 2023; Rieger et al., 2018). In the context of hate speech, counterspeech functions as a direct response—aiming not only to refute harmful content and mitigate its impact, but also to support targeted individuals (Rieger et al., 2018). Despite its promise, engaging in online counterspeech poses significant challenges; individuals often encounter barriers such as fear of retaliation and limited writing skills (Mun et al., 2024; Ping et al., 2024). Research indicates that those facing these challenges tend to be less satisfied with their counterspeech, experience greater difficulty in crafting responses, and doubt the overall effectiveness of their interventions (Ping et al., 2024).
Rapid advances in natural language processing (NLP) and large language models (LLMs) have prompted researchers to examine how AI can be leveraged to help users combat online hate (Yu et al., 2022; Zhu and Bhat, 2021; Saha et al., 2022; Garland et al., 2020). Researchers have explored various NLP applications—from hate speech classification (Yu et al., 2022; Garland et al., 2020) and counterspeech tone recognition (Bonaldi et al., 2024) to automated counterspeech generation (Zhu and Bhat, 2021; Saha et al., 2022). However, studies have revealed significant limitations in fully AI-generated counterspeech, including concerns about lack of personal voice and inability to convey authentic human emotions when addressing sensitive topics (Baider, 2023; Ping et al., 2024; Mun et al., 2024; Hangartner et al., 2021). This highlights the need for a more balanced approach—one that leverages AI’s capabilities while preserving the human voice and emotional authenticity of users in counterspeech creation.
To address this, we developed CounterQuill, an AI-mediated counterspeech writing partner that guides users through a three-step process:
  • (1) a learning session to help users better understand the impact and nature of online hate speech and counterspeech
  • (2) a brainstorming session that supports users in identifying key elements of hate speech and exploring counterspeech strategies
  • (3) a co-writing session that helps users draft and refine their counterspeech based on sessions (1) and (2).
The development of CounterQuill was informed by a preliminary survey (N=458) aimed at understanding user perspectives on the role of AI in counterspeech engagement. Survey results showed that while some users saw potential in AI tools like ChatGPT for writing counterspeech, others were concerned about preserving authenticity and personal expression when using AI for counterspeech writing.
In line with established research practices, we evaluated CounterQuill against ChatGPT—a baseline widely used in studies evaluating AI writing tools (Zhang et al., 2023; Reza et al., 2024; Lin et al., 2024). This comparison allowed us to assess whether CounterQuill’s distinctive features—its integrated learning and brainstorming components alongside its collaborative writing environment—provide meaningful advantages over standard AI assistance. Specifically, we examined whether these additional components help users craft counterspeech they perceive as both effective and authentic to their own voice.
To this end, we conducted a within-subjects study with 20 participants, randomly assigning the order of conditions (half starting with CounterQuill and half with ChatGPT) to minimize potential biases from sequencing effects (Ryan and Morgan, 2007). Our work addresses the following research questions:
  • RQ1. How does AI-mediated learning sessions shape users’ understanding of hate speech and counterspeech?
  • RQ2. How does AI-mediated brainstorming sessions affect users’ confidence and ability to identify key elements of hate speech and counterspeech strategies?
  • RQ3. How does engaging in collaborative co-writing with AI influence users’ self-efficacy in composing counterspeech and their perceptions of the co-authored response?
Contributions: The intellectual merits of this work lie in the design of an AI-mediated counterspeech system and the potential social impact of human-AI collaboration.
  • We show that CounterQuill assists users in identifying the core elements of hate speech, such as the targeted individual or group’s identity and the dehumanizing actions. The highlighting practice and AI-generated feedback enhance participants’ ability to recognize these key components.
  • We find that CounterQuill increases users’ confidence in crafting empathetic counterspeech. The system’s brainstorming session combined with the co-writing session that provides targeted suggestions and rewriting assistance, empowers users to compose empathetic counterspeech.
  • We uncover a relationship between user cognitive effort and behavioral outcomes in AI-mediated communication (AIMC) systems. Our findings suggest that when users invest more time and effort in collaborating with an AI system to write counterspeech, they develop a stronger sense of ownership over the co-authored counterspeech, increasing their likelihood of posting it on social media.
  • Finally, we offer key design implications for AI-assisted counterspeech writing that balance system guidance with user autonomy, prioritizing meaningful user engagement with AI to enhance both the quality of Counterspeech and users’ sense of ownership over the co-created content.
2. Related work
2.1. Online Hate and its Dangers
Online hate – or cyberhate – is a type of cyberviolence that uses computer technology to express hatred toward a group based on the group’s race, ethnicity, gender, gender identity, sexual orientation, national origin, religion, or other group characteristic (Perry, 2009; Räsänen et al., 2016). It differs from cyberbullying because of this collective focus. Like terrorism, online hate can advocate the hatred of one group or the hatred of numerous groups. Exposure to online hate has increased in the U.S. (Costello and Hawdon, 2020), with over 60 percent of Americans ages 18 to 25 being exposed to it in 2023 (Costello et al., 2023; Hawdon, 2024). But it is not limited to the U.S., it is a global phenomenon (Reichelmann et al., 2021; Commission, 2018). Although not everyone who sees online hate suffers negative consequences, exposure to it is correlated with both short-term and long-term emotional distress, such as mood swings, anger, feelings of isolation, decreased levels of trust, and fear (Costello and Hawdon, 2020; Foxman and Wolf, 2013; Näsi et al., 2015; Nyman and Provozin, 2019). Because they are still forming their identities, young people are especially vulnerable to negative effects of online hate (Billett, 2014). Given the amount of online hate that can be found online and the dangers it can pose, it is imperative that we find effective tools for detecting it and combatting it. Recent advances in automated detection have been made, but detecting it is not enough. We also need to find ways to lessen its impact. Combatting online hate with counterspeech is one such strategy.
2.2. Human-AI Interaction with Generative AI
Recent advancements in Generative AI like GPT-4, Bard, and Llama have enabled their integration into various application domains (Achiam et al., 2023; Rahsepar et al., 2023; Touvron et al., 2023), causing interest in how users interact, collaborate, and co-write with these AI systems (Wu et al., 2023; Tang et al., 2024; Wang et al., 2023; Duin et al., 2023; Balel, 2023; Lee et al., 2024a, 2022). In the field of co-writing with AI, researchers have shed light on aspects of user engagement, adaptation, and reliance on AI-generated content (Wu et al., 2023; Lewis and Moorkens, 2020; Lee et al., 2024a). These studies clarify activities ranging from using AI suggestions for conceptualization and content generation to evaluating (Desmond et al., 2024; Lee et al., 2024b), editing, and refining machine-generated text (Yang et al., 2022; Balel, 2023; Coenen et al., 2021; Zhang et al., 2023), such as creative writing (Coenen et al., 2021; Yang et al., 2022), technical documentation (Zhang et al., 2023; Duin et al., 2023), and online communication (Jakesch et al., 2023). Despite such advancements, there are limitations in AI-assisted writing, such as incorporating user instructions (Yeh et al., 2024; Wasi et al., 2024), handling biased AI-generated content (Jakesch et al., 2023), and navigating challenges related to AI’s understanding of user intent or desired tone (Ping et al., 2024; Lee et al., 2024a). For instance, Osone et al. (2021) identified challenges in collaborating with AI in creative writing, where experienced writers exhibited dissatisfaction with AI-generated content due to misalignment with expected storylines (Osone et al., 2021). Meanwhile, as Zhou et al. (2023) revealed that generative AI creates disinformation by both reshuffling human-made data and manipulating language to produce content that seems credible to targeted audiences.(Zhou et al., 2023).
As the integration of AI into writing platforms becomes more prevalent, not only writers and researchers but also social media developers and users will incorporate AI tools into their workflows to enhance the co-creation of social media content (Mun et al., 2024; Ping et al., 2024). Particularly, the collaboration between humans and AI in generating counterspeech involves a nuanced understanding of language, tone, and context to counteract harmful narratives online (Mun et al., 2024; Ping et al., 2024; Ray, 2023). For instance, Ray (2023) and Ping et al. (2024) highlighted that researchers, while crafting counterspeech, find AI tools beneficial in generating initial ideas. Still, the refinement of language, ensuring accuracy, and constructive tone often require human intervention (Ping et al., 2024; Ray, 2023). Additionally , the design and implementation of AI-assisted tools for counterspeech present challenges, such as ensuring that the AI-generated content involves sufficient human participation (Ping et al., 2024; Ray, 2023), aiding users in providing ample background and support when co-writing counterspeech with AI (Ping et al., 2024), and assisting users in maintaining the credibility and efficacy of the counter-argument when crafting counterspeech (Mun et al., 2024; Ping et al., 2024).
Overall, while previous research has provided valuable support for initial counterspeech creation, conscious human engagement is needed to ensure its accuracy and effectiveness in the Human&AI co-creation process (Mun et al., 2024; Ping et al., 2024). Our research aims to bridge the gap by exploring ways to enhance human-AI collaboration so that AI-generated counterspeech resonates effectively with the target audience and context.
2.3. Counterspeech Generation
Text generation, especially counterspeech generation, is one of the many open challenges in NLP research that has made breakthroughs in recent years(Zhu and Bhat, 2021; Chung et al., 2020, 2021; Saha et al., 2024). Researchers have explored various aspects of counterspeech generation using LLMs, such as generating contextually relevant responses (Hassan and Alikhani, 2023; Zhu and Bhat, 2021), knowledge-grounded counterspeech generation (Chung et al., 2021), and ensuring that the generated counterspeech adheres to ethical and societal norms (Mun et al., 2023). For instance, Chung et al. (2022) explored whether adding relevant external information could improve LLMs’ ability to generate counterspeech to hate speech (Chung et al., 2021). By extracting keyphrases from the hate speech and finding related outside knowledge, they were able to generate more suitable and informative counterspeech. Meanwhile, Hassan et al. (2023) have developed a novel framework based on theories of discourse to study the inferential links connecting counterspeech to hateful comments (Hassan and Alikhani, 2023). They proposed a taxonomy of counterspeech derived from discourse frameworks and discourse-informed prompting strategies to generate contextually-grounded counterspeech (Hassan and Alikhani, 2023).
Despite LLMs’ ability to generate counterspeech, we do not know whether machine-generated counterspeech aligns with people’s preferences or whether people would actually use it to counter hate online (Mun et al., 2024; Ping et al., 2024). Prior research shows that people are hesitant to use counterspeech entirely generated by AI because it strips away their personal voice and raises concerns about authenticity (Ping et al., 2024). On the other hand, some believe AI assistance in counterspeech-writing can reduce the emotional burden of responding to online hate and help them articulate their thoughts better (Ping et al., 2024). Researchers in HCI and CSS emphasize the importance of a human-centered approach in understanding the use of AI-generated content. Likewise, we argue that effectively using LLMs to counter hate speech in the real world requires attention to human involvement, preferences, and concerns. Our work adopts this approach to improve LLMs’ ability to generate human-centered counterspeech and to explore how humans can co-create counterspeech with AI in alignment with their preferences, personal voices, and ideas.
2.4. Ownership in Human-AI Collaboration
Previous research has demonstrated that ownership extends beyond physical or tangible objects, encompassing everything from people to text to ideas (Heider, 1958; Mieczkowski, 2022). Regarding how interaction with AI systems affects the human sense of ownership, past studies have found that the more people engage in their interactions with AI, the more likely they are to develop a strong sense of ownership over the outcomes of their co-creation with the AI (Louie et al., 2021; Pierce et al., 2003; Louie et al., 2020). For instance, in Louie et al.’s (2021) study of AI music co-creation, participants indicated that they felt the final product did not belong to them due to their minimal control over the AI-generated suggestions (Louie et al., 2021). Furthermore, Draxler et al.’s study of AI-assisted writing suggested that users’ perceived control over AI-generated content and their feeling of leadership in the interaction significantly impact their sense of ownership (Draxler et al., 2024).
As AI-assisted writing continues to evolve, people are increasingly using with AI to create content on social media; one such example is counterspeech writing (Buerger, 2021; Saha et al., 2022; Mun et al., 2024). While previous research has shown that AI can help people overcome some writing barriers in the process of creating counterspeech (Mun et al., 2024; Zhu and Bhat, 2021; Saha et al., 2024), studies by Kumar (2024) and Ping et al. (2024) have highlighted that people feel AI, such as ChatGPT, prevented users from feeling a sense of ownership over the words and ideas expressed (Kumar, 2024; Ping et al., 2024). This loss of ownership has also raised concerns about authenticity and ethics (Kumar, 2024), which may limit their willingness to use the generated counterspeech.
Overall, the growing collaboration between humans and AI in content creation has led to an interplay of user effort (Mieczkowski, 2022; Louie et al., 2021), perceived ownership (Mieczkowski, 2022; Kumar, 2024; Louie et al., 2020; Pierce et al., 2003), and ethical considerations (Kumar, 2024; Ping et al., 2024). Our research aims to further explore the relationship between people’s effort, perceived ownership, and future behavioral outcomes in human-AI collaboration.
3. DESIGN GOALS
We adopt a human-centered approach in designing CounterQuill as an interactive human-AI system that empowers users with relevant knowledge (DG1), ideation space (DG2), and collaborative writing assistance (DG3) rather than simply generating counterspeech. By doing so, our system aims to help users engage in the process of crafting effective and empathetic counterspeech that incorporates their preferences and thought process in countering online hate, rather than solely delivering a machine-generated output. Our system is designed to achieve the following design goals:
  • DG1. Help users understand hate speech and counterspeech and their impact.
  • DG2. Help users ideate counterspeech ideas through a brainstorming process.
  • DG3. Help users draft and refine their counterspeech with AI assistance.
3.1. Help Users Learn About Hate Speech & Counterspeech and Their Impact (DG1)
Studies show many internet users, especially young people, encounter hate speech regularly but struggle to identify or report it (Räsänen et al., 2016; Anti-Defamation League, 2019). A survey found 65% of American youth had seen online hate speech, yet only half recognized it as harmful—highlighting a critical need for better education on identifying these behaviors (Anti-Defamation League, 2019). To address this issue, organizations have launched campaigns to teach people how to recognize hate speech, but challenges remain due to its complex nature. Hate speech isn’t always obvious. While some attacks use direct slurs or threats (explicit hate speech) (Reichelmann et al., 2021; Howard, 2021), others hide behind sarcasm (Ocampo et al., 2023), coded language (ElSherief et al., 2021; Ocampo et al., 2023), or subtle stereotypes (implicit hate speech) (Gunturi et al., 2024). The subtleties of implicit hate speech make it hard for people to detect, as meaning shifts based on intent, audience, and cultural cues (Gunturi et al., 2024).
Compounding this, if counterspeech is not delivered carefully, it can often escalate tensions in online hate speech discussions (Howard, 2021), as aggressive or lecturing responses tend to provoke further hostility (Howard, 2021; Chung et al., 2023). To address this challenge, researchers have explored more effective intervention strategies. For example, Hangartner et al. show that empathy-based counterspeech - humanizes the victim and reminds the perpetrator that people can be hurt by their behavior - is more likely to lead to apologies or deletion of hateful posts by bad actors (Hangartner et al., 2021). However, most people lack awareness of these counterspeech strategies and their relative effectiveness (Ping et al., 2024). To address this knowledge gap, we developed our learning session aiming to educate users about both hate speech and counterspeech and understand the impact of different response approaches.
3.2. Help Users Ideate Counterspeech Ideas Through a Brainstorming Process (DG2)
Prior studies have shown that many people hesitate to engage in online counterspeech due to the cognitive load (Mun et al., 2024; Buerger, 2021; Ping et al., 2024). This arises from the difficulty in identifying the key hate elements of hate speech, such as aspects of an individual or group’s identity and the dehumanization of that identity (Waltman, 2018; Harel et al., 2020) - a task that is particularly challenging in implicit hate speech. Furthermore, many people may lack the skills to counter hate speech (Ping et al., 2024). For instance, a study by Costello et al. (2017) found that although participants (N = 963) encountered online hate speech, only 3.1% indicated they would respond with a counterspeech, while others were unsure how to respond effectively(Costello et al., 2017).
To help users overcome such barriers and practice counterspeech skills, we propose a brainstorming approach to guide counterspeech ideation process. We adopted the traditional brainstorming approach (TBS) (Gallupe et al., 1991; Yu-Han and Chun-Ching, 2023) and extended it by drawing on prior HCI studies on AI technologies that facilitate brainstorming (Lavrič and Škraba, 2023; Zhang et al., 2023; Yu-Han and Chun-Ching, 2023; Cepolina et al., 2022). For example, researchers in their studies use LLMs to help people generate (Zhang et al., 2023; Yu-Han and Chun-Ching, 2023), refine (Yu-Han and Chun-Ching, 2023; Geerts et al., 2021), and evaluate ideas (Cepolina et al., 2022). Zhang et al. created Visar, a human-AI collaborative writing assistant that helps writers brainstorm and revise hierarchical goals within their argumentative writing context (Zhang et al., 2023). In addition, Lubart (2005) and Chiu et al. (2023) (Lubart, 2005; Yu-Han and Chun-Ching, 2023), identified four key components in HCI brainstorming: (1) AI engaging in communication and guidance with humans during the thinking process; (2) AI providing targeted feedback; (3) AI and humans collaborating to organize brainstorming results; and (4) AI encouraging reflective behavior by integrating human-AI collaboration into the creative process. Built on previous research, we developed our brainstorming session aiming to help users ideate counterspeech strategy.
3.3. Help Users Draft and Refine Their Counterspeech With AI Assistance (DG3)
Even after learning about hate speech and ideating potential counterspeech strategies, many individuals still face barriers when actually writing counterspeech. Recent research identifies several key challenges in writing counterspeech, including emotional burden, time demands, and perceived lack of expertise in creating impactful responses (Ping et al., 2024).
To address these challenges, researchers have explored the potential of LLMs and AI technologies to support counterspeech writing. However, while AI assistance offers some advantages—such as reducing cognitive load and enhancing rhetorical effectiveness—are users willing to embrace this technology? To better understand people’s perspectives on AI-assisted counterspeech, we conducted a preliminary survey (N=458) before designing DG3 (Appendix A). The survey explored three key questions:
  • (1)
    Have you ever used an AI language model such as ChatGPT? (0 = No, 1 = Yes)
  • (2)
    If you were writing a counterspeech on social media, would you use AI technology like ChatGPT to help you write it? (1 = Definitely not, 5 = Definitely yes)
  • (3)
    Why would or wouldn’t you use AI technology to help you write a counterspeech?
Our results indicate that among participants with prior experience using ChatGPT (N=162), the mean likelihood of using AI for counterspeech was 2.46 (SD = 1.17), which suggests while some participants were open to trying ChatGPT, they also had some reservations. To understand the reasons behind this, we conducted a qualitative analysis on their open-ended responses, which revealed four key themes (Table 1).
Table 1. Positive Perspectives and Concerns: Participant Feedback on Using ChatGPT for Counterspeech
Theme Illustrative Quotes
Positive Perspectives Efficiency and Convenience ”I think [ChatGPT] is better and faster at putting together coherent sentences that get my point across than I am.” ”[ChatGPT] can give me ideas quickly, and I can elaborate with my own perspective of the facts.”
Reduced Emotional Burden ”I would probably have the [ChatGPT] help, partially because I just don’t have the energy for that sort of thing anymore.” ”I feel like [ChatGPT] would be able to refute it with facts and logic better than me because counterspeech is an emotional burden, and I get overwhelmed.”
Concerns and Limitations Authenticity and Ethical Concerns ”It’s not my voice. It’s not my perspective. Personally, I’d be ashamed to utilize artificial intelligence for counterspeech.” ”Because then it wouldn’t even be MY counterspeech. Why would I use AI to write MY opinion? It’s stupid.” ”I would want to build the skills to effectively and reliably write such speech myself.”
Lack of Emotional, Human, or Personal Touch ”I will write counterspeech better since I can incorporate personal experiences to help the other person better understand my true intentions.” ”It can use logic, but not empathy, to write counterspeech.” ”I would rather tailor my response to be exactly what I’m thinking. I’m not sure it could fully express what I want to say, and it may lack nuance.”
As illustrated in the themes in (Table 1), participants recognized both benefits and limitations of AI assistance in counterspeech writing. While they appreciated AI’s potential to improve efficiency and reduce emotional burden, they expressed concerns about maintaining authenticity and personal voice in their responses. ChatGPT consistently emerged as participants’ primary reference point when discussing AI assistance, with many explicitly mentioning it as representative of current AI capabilities. This widespread familiarity with ChatGPT among participants, combined with established research practices in prior work evaluating AI writing tools (Zhang et al., 2023; Reza et al., 2024; Lin et al., 2024), provided a clear rationale for selecting ChatGPT as our baseline for comparison.
4. System Implementation
CounterQuill was developed using React for the front end, and Python, OpenAI, and MySQL for the back end. The interface used Tailwind CSS with minimal aesthetics to indicate the system features. The prompt content used in CounterQuill is included in Appendix F.
4.1. CounterQuill Learning Session
The learning session has three main sections: 1 What is hate speech 2 Learning about counterspeech and 3 Testing your understanding (Figure 2).
Figure 2. User Interface for DG1 (Help people understand hate speech & counter speech and their impact). (1) Section for educating people to understand hatespeech; (2) Section for educating people to understand counterspeech; (3) Test people’s understanding of hatespeech & counterspeech.
The content for the Learning session was developed through an iterative design process informed by literature and expertise from two sociologists, and two HCI experts:
1What is hate speech
: it contains three learning content subsections, which users can navigate through by clicking the Next button to progress and the Previous button to review. It covers three crucial topics for comprehending hate speech: (1) The definition of hate speech; (2) The explicit or implicit hate speech, and (3) How does hate speech affect us?
2Learning about counterspeech
: it has three subsections of learning content. It focuses on three key aspects of counterspeech: (1) What is counterspeech? (2) What makes counterspeech effective? and (3) What is an empathy-based counterspeech?
3Testing your understanding
: it includes four questions to assess users’ comprehension of the learning material covered in the previous sections. These questions evaluate users’ understanding of (1) the definition of hate speech, (2) the effects of explicit and implicit hate speech, (3) the definition of counterspeech, and (4) the concept of empathy-based counterspeech (Appendix E).
4.2. CounterQuill Brainstorming Session
To help users ideate counterspeech ideas, we designed the Brainstorming session with two key sections: (A) a section for users to highlight key elements of hate speech (Figure 3) and (B) a section to prompt users to reflect on the hateful statement and emotional impact on the targeted individuals (Figure 4). Section (A) includes: 1 Tutorial on highlighting practice and 2 Highlighing area. Section (B) includes: 3 Brainstorming through Q&A and 4 CounterQuill suggestions on brainstorming results.
Figure 3. User Interface for DG2 (Help users ideate counterspeech ideas through a brainstorming process)’s Highlighting Practice: (1) Tutorial on highlighting practice; (2) Highlighting area: identify the individual/ group’s identity and dehumanizing actions.
4.2.1. Highlighting Practice
1Tutorial on highlighting practice
: it features a Brainstorming Guidance Agent (BGA) that communicates the brainstorming tasks and steps. BGA demonstrates how to highlight aspects of an individual or group’s identity in yellow and actions that dehumanize them in green.
2Highligting area
: users manually select and highlight the identity aspects in yellow and the dehumanizing actions in green. Once users have completed highlighting the content, they can click the ”Done” button, triggering Brainstorming Feedback Agent (BFA) to semantically compare the user-highlighted content with the system’s predefined correct answer (Luo et al., 2023).
For example, in the hate speech statement “I saw a random black man jogging nearby our house today and now I feel unsafe walking around my own neighborhood,” the correct semantic understanding of the dehumanizing action could be “feel unsafe”. If the user selects “unsafe” or “I feel unsafe”, they have correctly understood the dehumanizing action. However, if the user selects “jogging nearby”, it is an incorrect understanding of the dehumanizing action, and the BFA will generate corresponding corrective feedback. Additionally, users can click the “View” button to see the differences between their understanding of identity and dehumanizing action compared to the system’s interpretation.
4.2.2. Question Answering
3Brainstorming through Q&A
: we designed two thought-provoking questions that encourage users to provide thoughtful written responses (Figure 4). Question 1 asks, “What negative stereotypes or assumptions about the targeted group or individual are suggested by the statement?” This question aims to let users think of harmful assumptions implied about the targeted group or individual. Question 2 asks, “Consider the feelings and experiences of someone who identifies with the group mentioned in the statement. How might this comment affect their sense of safety, belonging, or self-esteem?” This question encourages users to reflect on the impact of identity-based dehumanization.
Figure 4. User Interface for DG2’s Question Answering: (3) Brainstorming through Q&A, (4) CounterQuill suggestion on brainstorming results.
4CounterQuill suggestions on brainstorming results
: after completing their input for both questions, users can click the “Done” button, prompting CounterQuill to provide suggestions based on the user’s responses. These suggestions include (1) feedback on the user’s understanding and (2) suggestions on strategies for crafting their counterspeech.
Users can then review CounterQuill’s suggestions and select the parts they find noteworthy for drafting their counterspeech. By selecting the relevant text and clicking “Take Notes” to highlight them, they create notes that CounterQuill stores in the backend database.
4.3. CounterQuill Co-Writing Session
To help users draft and refine their counterspeech, we designed a co-writing session with CounterQuill. This session includes four main sections: 1 Tutorial about writing assistant, 2 Brainstorming notes section for users to review, 3 Text area for users to write counterspeech, and 4 CounterQuill writing assistance tool.
1Tutorial about writing assistant
: the Writing Guiding Agent (WGA) interacts with users to introduce the writing tasks, including the hate speech they need to counter and a video tutorial on co-writing with CounterQuill. The tutorial video demonstrates how users can obtain writing suggestions and refine their counterspeech with CounterQuill.
2Brainstorming notes section
: it displays the notes that users have taken by clicking the “Take Notes” button during the brainstorming session. These notes are the ones users find noteworthy and helpful for writing their counterspeech. To implement this feature, the WGA retrieves user-selected content from our backend database (see § 4.2.2) and displays their notes to users.
Figure 5. User Interface for DG3 (Help users draft and refine their counterspeech with AI assistance). (1) Tutorial about CounterQuill writing assistant, (2) Brainstorming notes taken by the users, (3) Text area for users to write counterspeech.
3Counterspeech writing area
: our system uses the user’s reflections on hate speech from the brainstorming session (i.e., their responses to the two questions considering the negative assumptions suggested by the statement and the feelings of those targeted) as an initial draft for their counterspeech. Users can then make initial additions, deletions, and modifications to the content based on their brainstorming notes. As users begin inputting content, our system automatically saves the input to the backend database. Once users complete their initial draft, they can select a portion or all of the text and click on the writing assistant button to invoke the help of the CounterQuill Writing Assistant (Figure 6 A ).
Figure 6. User Interface for DG3 - Select text and use CounterQuill writing assistant to refine the content. (A) CounterQuill Writing Help Button, (B) CounterQuill writing assistant’s features, (C) Sample result polished by CounterQuill
4CounterQuill writing assistant
: we have designed features for grammar checking, polishing content to an empathetic tone, and rewriting content using the user’s brainstorming notes 1 and 2. Finally, to help users customize their writing, we have designed a “Customize Your Writing” area where users can input their desired final writing goals to align with their preferences. Specifically, when users click on the corresponding button (Figure 6 B ), CounterQuill will rewrite the text according to the user’s needs (see Appendix F for detailed prompts) and present the result in (Figure 6 C ). Finally, if users approve of the CounterQuill’s output, they click “Insert” to add it to the editing area; otherwise, clicking “Retry” generates a new rewrite.
5. User Study
We conducted a user study with 20 participants to evaluate CounterQuill. The study was approved by an Institutional Review Board (IRB), and designed to take 60 minutes for each participant.
Table 2. Demographic information of CounterQuill user study participants.
ID Gender Ethnicity Age Education
P1 Male Asian/Asian American 21 Bachelor’s degree
P2 Male Black/African American 26 Some college
P3 Female Black/African American 24 Associate degree
P4 Female White/European American 32 Associate degree
P5 Female White/European American 44 Bachelor’s degree
P6 Female Asian/Asian American 32 Bachelor’s degree
P7 Male Asian/Asian American 23 Bachelor’s degree
P8 Male Hispanic/Latina/o/x 21 Bachelor’s degree
P9 Male Asian/Asian American 23 Bachelor’s degree
P10 Male Middle Eastern/North African 24 Doctorate degree
P11 Female White/European American 24 Master degree
P12 Female White/European American 26 Master degree
P13 Male Black/African American 34 Bachelor’s degree
P14 Female Black/African American 29 Bachelor’s degree
P15 Female Asian/Asian American 27 Master degree
P16 Male White/European American 32 Master degree
P17 Male Middle Eastern/North African 21 Bachelor’s degree
P18 Male White/European American 22 Bachelor’s degree
P19 Male White/European American 31 Bachelor’s degree
P20 Female Black/African American 34 Associate degree
5.1. Participants
We recruited 20 native English speakers (9F, 11M, ages 20-44, mean = 26, SD = 5.7) through word-of-mouth, LinkedIn, Reddit, and X (Table 2). Five participated in-person at a usability lab and fifteen joined virtually via Zoom. Participants consented to the use of their survey responses and interview recordings for data analysis and received \$20 gift cards as compensation. There were two eligibility requirements to participate: prior experience with generative AI tools (e.g., ChatGPT, Claude) and previous encounters with hate speech where they had engaged in counterspeech. All 20 participants have used AI writing tools and have responded to online hate speech with a counterspeech. Six had previously used ChatGPT to write online counterspeech in the past.
5.2. Study Protocol
Each 60-minute study session followed a within-subjects design. During sessions, participants wrote counterspeech using ChatGPT (20 minutes) and co-wrote with CounterQuill (40 minutes). Before the study, participants received a pre-study email with a tutorial video explaining CounterQuill’s features.
5.2.1. Conditions
We implemented two distinct conditions to evaluate the AI-assisted counterspeech writing, drawing on our findings in a preliminary survey (3.3) and previous research using GPT-based systems (Zhang et al., 2023; Reza et al., 2024; Lin et al., 2024) such as GPT-3.5 as baseline:
  • ChatGPT (Baseline): Participants used ChatGPT to write counterspeech for the hateful content. ChatGPT allowed users to input prompts and receive generated responses from the GPT-3.5-turbo language model, which is the same model used in CounterQuill.
  • CounterQuill: Participants used CounterQuill, which includes a learning session (help users understand hate speech and counterspeech), a brainstorming session (step-by-step counterspeech ideation support), and a writing session (co-writing with CounterQuill for counterspeech).
To minimize potential biases, we used a Latin Square experimental design (Ryan and Morgan, 2007) to ensure a balanced sequence of task topics and conditions for each study session: with half the participants starting with our system and the other half beginning with ChatGPT. To assess the potential impact of the system usage order on participants’ perceptions and the overall study outcomes, we conducted Welch’s Two-Sample t-tests that accounts for unequal variances between groups (Sakai, 2016), as detailed in Tables A1 and A2. The results confirmed that the sequence—whether participants began with ChatGPT or CounterQuill—did not significantly influence their performance or feedback on the system.
5.3. Study Procedure
At the beginning of each study session, the experimenters collected informed consent from the participants. For the user study, each participant was presented with 20 hate speech examples and asked to select one of them. The hate speech data was sourced from Ping’s (2024) research (Ping et al., 2024), which included a total of 400 hate speech instances evenly distributed across five hate speech themes: race, gender, sexual orientation, disability, and religion. To ensure a balanced representation of themes, each participant was randomly assigned 20 hate speech examples, with an average of 4 examples per theme. Appendix H lists the hate speech instances used by each of the 20 participants.
In the baseline condition, participants used ChatGPT to compose counterspeech responses to their selected hateful content, then completed a post-study questionnaire featuring NASA-TLX and 6 questions rating their counterspeech writing experience (Appendix D). In the CounterQuill condition, participants underwent a learning session about hate speech and counterspeech concepts, engaged in a structured brainstorming session for step-by-step ideation, and used the CounterQuill co-writing assistant before completing the same questionnaire. Both conditions concluded with a 20-minute semi-structured interview.
6. Result
6.1. Quantitative Evaluation
6.1.1. Comparative NASA-TLX Analysis
Table 3 summarizes the ratings for each dimension of the NASA Task Load Index (TLX) across two conditions: using ChatGPT (baseline) and using CounterQuill. We evaluated the differences between these conditions on each dimension using paired-sample $t$-tests (Mowery, 2011). To further contextualize these findings, Table 4 provides a comparison of the average time spent on counterspeech composition in each condition.
Table 3. Comparison of NASA-TLX scores between ChatGPT and CounterQuill for counterspeech writing tasks. The results of paired $t$-test show statistically significant differences in five out of six TLX subscales, with CounterQuill scoring higher in Physical Demand, Temporal Demand, Performance, and Effort, while ChatGPT scored higher in Frustration.
[…]
NASA-TLX Items $Mean\pm SD$ $t$-value $p$-value 95% CI
ChatGPT CounterQuill
Mental Demand 4.53$\pm$1.70 4.94$\pm$1.39 -1.237 0.234 [-1.117, 0.294]
Physical Demand 2.06$\pm$1.52 3.12$\pm$1.73 -3.816 0.002** [-1.647, -0.471]
Temporal Demand 2.65$\pm$1.32 4.47$\pm$1.46 -7.410 ¡0.001*** [-2.345, -1.302]
Performance 3.88$\pm$1.17 6.35$\pm$0.70 -7.177 ¡0.001*** [-3.200, -1.741]
Effort 2.53$\pm$1.07 3.82$\pm$1.67 -3.235 0.005** [-2.142, -0.446]
Frustration 3.06$\pm$1.95 1.47$\pm$0.72 3.628 0.002** [0.660, 2.516]
The analysis of the NASA-TLX dimensions revealed several notable differences between the two conditions. While Mental Demand was slightly higher with CounterQuill (though not statistically significant), Physical Demand, Temporal Demand, and Effort were higher when using CounterQuill. Performance was also significantly better with CounterQuill, showing a large effect size. Conversely, Frustration was the only dimension that was significantly lower when using CounterQuill compared to ChatGPT.
Table 4. Comparison of Average Time Spent on Counterspeech Composition
Task Average Time (min)
CounterQuill Learning 8.72 $\pm$ 1.02
CounterQuill Brainstorming 10.32 $\pm$ 1.55
CounterQuill Writing 10.48 $\pm$ 1.71
ChatGPT Writing 9.11 $\pm$ 1.06
In summary, CounterQuill demands greater effort, time, and energy than ChatGPT, but delivers higher perceived performance outcomes (as measured by NASA TLX ’Performance’), with participants reporting a 63.66% improvement. This stems not only from participants’ satisfaction with their co-written counterspeech but also from their strong performance in the post-learning session quiz, where 15 of 20 participants achieved perfect scores (91.25% average accuracy). Furthermore, CounterQuill reduced participants’ frustration levels by 51.96% compared to ChatGPT (as measured by ’Frustration’).
6.1.2. User Experience Ratings
Our study further investigated participants’ experiences and perceptions when using ChatGPT and CounterQuill for counterspeech writing. To gain deeper insights, we designed 6 custom questions (see Appendix D for detailed post-study questionnaire) that focused on specific aspects of the participants’ experience, such as confidence in identifying hate speech, effectiveness of the tools in providing guidance, and satisfaction with the final counter-speech. All questions were rated on a 7-point Likert scale.
[…]
Category of Question $Mean\pm SD$ $t$-value $p$-value 95% CI
ChatGPT CounterQuill
Confidence in Identifying Hate Speech 4.18$\pm$1.88 6.24$\pm$0.75 -4.280 ¡0.001*** [-3.079, -1.039]
Effectiveness in Brainstorming Ideas 3.24$\pm$1.44 6.35$\pm$0.70 -8.849 ¡0.001*** [-3.865, -2.371]
Confidence in Writing Counterspeech 3.29$\pm$1.83 6.18$\pm$0.81 -7.356 ¡0.001*** [-3.713, -2.052]
Engagement & Motivation While Using AI 3.59$\pm$1.73 6.00$\pm$0.87 -5.734 ¡0.001*** [-3.303, -1.520]
Satisfaction with the Final Counterspeech 4.00$\pm$1.58 6.47$\pm$0.51 -7.668 ¡0.001*** [-3.154, -1.788]
Willingness to Share Counterspeech Online 3.53$\pm$1.46 6.29$\pm$0.77 -10.444 ¡0.001*** [-3.326, -2.204]
Overall, CounterQuill demonstrated statistically superior outcomes compared to ChatGPT ($p$¡0.001). As shown in Table 5, key findings include:
  • Learning Session: Participants gained higher confidence in identifying online hate speech (HS Identification Confidence: 6.24 vs. 4.18).
  • Brainstorming Session: Participants rated CounterQuill as more effective for generating ideas (Brainstorming Effectiveness: 6.35 vs. 3.24) and reported stronger self-efficacy in ideating counterspeech (Self-Efficacy: 6.18 vs. 3.29).
  • Co-Writing Session: The collaborative process led to greater satisfaction with the final counterspeech (Satisfaction: 6.47 vs. 4.00) and increased willingness to share it online (Willingness to Post: 6.29 vs. 3.53).
6.2. Insights from the Learning Session
6.2.1. Learning About Online Hate Speech and Their Impact
The learning session enhanced participants’ ability to distinguish between explicit and implicit hate speech. While explicit hate speech was widely recognized, many participants (19 — P1–P13, P15–P20) initially struggled with identifying implicit hate speech. P7 noted, “Before the session, I knew explicit hate speech was bad, but I didn’t really think this (implicit hate speech) could also be hateful.”
After the session, participants (15 — P1–P5, P7, P8, P10–P13, P15–P20) expressed surprise at how subtle yet harmful implicit hate speech can be. P13 observed, “After finishing the learning session, I feel that because implicit hate speech is subtle, it requires deeper reflection to identify.” Similarly, P12 emphasized how implicit hate speech is often normalized in everyday discourse:
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=artifact, linewidth=0pt, backgroundcolor=artifact!10 {mdframed}
Participant 12: Now, I think implicit is harder to recognize because many people are not aware that what they’re saying is actually hate speech. They might think it’s just a normal thing to say, but there’s implicit[ly] hateful meaning behind it. So, implicit hate speech can be trickier to point out because it’s more subtle and social[ly] normalized.
6.2.2. Understanding Counterspeech and Their Impact
Before the session, participants had varying levels of familiarity with counterspeech. Some (14 — P1–P4, P8, P10–P13, P14–P17, P19) understood it as a general response to hate speech, while others (6 — P5, P6, P7, P9, P18, P20) were less familiar with the concept.
After completing the session, participants (18 — P2–P12, P14–P20) described counterspeech as a respectful and non-escalatory way to challenge hate speech. P4 provided a clear example:
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=artifact, linewidth=0pt, backgroundcolor=artifact!10 {mdframed}
Participant 4: I did respond to some hate [speech] before, but I never knew the term (of that behavior) … Counterspeech is a respectful way to address hate speech. It addresses the targets [as] human, and we should treat each other kindly … It’s a way to respond to hate [without] more hate, but get some understanding.
6.2.3. Discovering Empathy-Based Responses
Few participants (4 — P2, P4, P8, P16) were previously aware of empathy-based counterspeech, but after the session, many (19 — P1–P12, P14–P20) recognized its effectiveness in de-escalating conflicts and fostering reflection.
[…]
Participant 6: I like what I learned empathy-based counterspeech (from the CounterQuill). If someone posted a hateful comment about gay people, I could respond like: “Hey, just so you know — my son is gay, and he’s a lot more than what you said in your comment. He has dreams, hopes, and feelings.” By sharing my story, I can encourage people to reconsider their views.
[…]
Participant 18: After the learning session, I had more to think about. … humorous counterspeech can still come across as implicitly hateful. For example, if someone posts a negative comment about body image and you respond with something like, “Well, you look ugly too,” it doesn’t really work. It’s just you’re joking back at them in a way of negativity. In my opinion, the best way to convey counterspeech is through empathy.
Overall, the learning session deepened participants’ understanding of hate speech and counterspeech, equipping them with more effective strategies for responding constructively.
6.3. Insights from the Brainstorming Session
6.3.1. Interactive Highlighting of Hate Speech Elements
Our brainstorming session introduced participants to a color-coded highlighting practice, helping them identify the key components of hate speech: (1) Aspects of an individual or a group’s identity; (2) Action or perception that dehumanizes the individual or group based on (1).
Prior to the highlighting practice, most participants (19 — P1 - 13, P15 - P20) had difficulty identifying specific components of hate speech. As P11 mentioned, “I knew hate speech targeted certain groups …, but I couldn’t really break it down further than that”. After completing the practice, participants (17 — P1–P13, P16, P17, P19, P20) reported that using different highlight colors helped them visually distinguish key components of hate speech, such as targeted identities and dehumanizing actions. P5 described the process:
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=expert, linewidth=0pt, backgroundcolor=expert!10 {mdframed}
Participant 5: The highlighting made it easy to get the main points [in the hate speech]. When I read the example, I got what was being said and the negative things [that targeted the group]. It showed what the hate speech was about and the specific bad things.
Many participants (15 — P1–P7, P9–P12, P15, P17, P19, P20) also found the system’s feedback on the users’ highlighted text using a side-by-side comparison of the correct highlights to be helpful. P13 explained how the feedback improved her awareness:
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=expert, linewidth=0pt, backgroundcolor=expert!10 {mdframed}
Participant 13: After comparing my highlights with the system’s suggestions, … I see them as a good reference to better understand [the hatespeech]. It gave me a sense of additional aspects to consider, like the focus on women [as the target identity], not just independent women, and how even mentioning instinct is part of the hate speech. Those were elements I didn’t recognize.
For reference hate speech: “All of the independent women I knew eventually settled down and got married. Your instinct will kick in soon.”
Similiarly, P6 summarized how the highlighting practice boosted confidence in identifying hate speech:
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=expert, linewidth=0pt, backgroundcolor=expert!10 {mdframed}
Participant 6: … Doing this highlight practice has really increased my confidence when it comes to finding targeted groups and dehumanizing language [in hate speech]. Before I tried this, I probably missed some of the more subtle stuff … Everyone knows hate speech is bad, but after highlighting it will let you understand why it’s so toxic.
6.3.2. Stimulating Reflective Ideation Through Guided Questions
Before engaging in brainstorming process, participants’ initial reactions to hate speech varied. Some (8 — P3, P5, P8, P11, P13, P14, P18, P19) mentioned that they would typically respond with anger or frustration, while others (6 — P1, P4, P6, P7, P10, P17) expressed uncertainty about how to counter hate speech effectively. However, many participants (16 —P1, P2, P4, P6, P7 - P11, P13, P14 - P18, P20) noted that answering the questions in the Brainstorming session led them to consider a more empathetic tone. For example, P8 commented:
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=expert, linewidth=0pt, backgroundcolor=expert!10 {mdframed}
Participant 8: When I was answering the first question, I was thinking about why this was wrong and trying to explain it. In the second part, I tried to show some empathy. So it changed how I approached it. If I didn’t answer those questions, I might have been more offensive in my response. But now I feel like I can write it more constructively, and say why it’s wrong and how the targeted people might feel.
Additionally, as mentioned by some participants (6 — P2, P7 - P9, P16, P20), the question-answering process guided them to consider the harmful impact and feelings of the targeted group. P2 elaborated on this shift in perspective:
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=expert, linewidth=0pt, backgroundcolor=expert!10 {mdframed}
Participant 2: … In the past when responding to or countering, I wasn’t asking myself questions like this internally, like what are the negative stereotypes, or considering the feelings of the targeted person and how it impacts them. It’s not something I intuitively think about. So I think this really leads me engage in that empathy-based counterspeech we talked about before. It brings my perspective on what’s happening, why it’s harmful, and what the impact is.
In addition to the questions-answering, most participants (16 — P1-P7, P9-P12, P15, P17, P19, P20) described benefiting from the system’s feedback on their response reflection and suggested counterspeech strategies (§4.2.2). These features helped reduce their anxiety and mental load, particularly when they were unsure how to begin writing counterspeech. P5 specifically noted that the system’s suggestions alleviated the stress of starting counterspeech from scratch.
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=expert, linewidth=0pt, backgroundcolor=expert!10 {mdframed}
Participant 5: After I got the feedback, there were things I wanted to add to my comments or input. It was like, ‘Oh, I could add that.’ The feedback gave me a good starting point for writing counterspeech. If you asked me to write a counterspeech from scratch, I would start from a place of uncertainty, and it might not be very straightforward. But the feedback gave me a clear target to move on.
Some participants (12 — P1-P6, P9-P12, P19, P20) noted that the system’s feedback was essential in helping them think in-depth about the key elements of the hate speech they were analyzing, which guided their approach to crafting a response. P20, for example, detailed:
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=expert, linewidth=0pt, backgroundcolor=expert!10 {mdframed}
Participant 20: I think the feedback is good because, firstly, the feedback seems to respond to things that I said. So, they identify points that are useful for writing counterspeech. But then later, it also expands more on [either] why what I said is relevant [or] how it can be used in counterspeech, and how that can lead to different kinds of responses to counterspeech. So, I think the format and the personalized response [by the system] are helpful. Yeah, it’s very detailed and valuable.
6.4. Insights from the Co-Writing Session
6.4.1. Transforming Brainstormed Ideas into Counterspeech
Most participants commented on the intuitive design and user-friendliness of CounterQuill’s writing interface. “The friendly layout and convenient features allowed participants to focus on composing their counterspeech” (P2), without being hindered by complex technology.
[…]
Participant 5: Before using CounterQuill, I wasn’t sure where to start when writing counterspeech. I had ideas, but I didn’t know how to structure them or make them sound convincing.
However, after using CounterQuill, participants reported that the integration of brainstorming notes into the writing interface provided a smooth transition from the ideation stage to the actual writing process. Having the ability to refer to these notes during writing was beneficial (11 — P2, P5, P6, P10, P11, P12, P14, P17 - P20), as it helped participants structure their responses and include key points. P18 elaborated:
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=training, linewidth=0pt, backgroundcolor=training!10 {mdframed}
Participant 18: … looking back at the notes was definitely helpful. I tried to structure my entire response based on the notes and the main points I had taken in the earlier sessions. Breaking things down in the notes beforehand really did help me in forming the text for my counterspeech … It just made the process a lot smoother.
Moreover, using the participants’ responses from the brainstorming session as an initial draft for the counterspeech reduced their psychological pressure and cognitive load of starting from scratch (14 — P2 - P6, P9, P11, P12, P14, P15, P17 - P20). This approach also improved participants’ confidence and helped them effectively capture the key points they wanted to address (5 — P2, P8, P10, P11, P12). P8 noted how this feature significantly increased their confidence.
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=training, linewidth=0pt, backgroundcolor=training!10 {mdframed}
Participant 8: … Starting a draft based on my own ideas or [responses] made the writing [process] less stressful. I felt like I had an initial point, and it increased my confidence [for writing counterspeech] and now I know where I need to start about the counterspeech.
6.4.2. Achieving a Sense of Ownership Through Collaborative Revision
Once participants completed the draft of their counterspeech, they could select specific content and request writing assistance from CounterQuill. Most participants (16 — P1, P3, P5 - P11, P14, P15 - P20) found that the fine-grained control over choosing particular sentences or paragraphs for AI-assisted polishing was more personalized than current conversational AI (i.e., ChatGPT), providing a level of customization that other tools don’t offer. For example, P1 emphasized this advantage:
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=training, linewidth=0pt, backgroundcolor=training!10 {mdframed}
Participant 1: I think the ability to select specific sentences for rewriting is valuable. I wrote a pretty good draft and I want to keep most of it, but maybe there are one or two sentences that I want to include and I don’t really know how to phrase them. Having this system in place is helpful because you can keep what you wrote originally while also editing it part by part with the AI, all within the context of the full response.
[…]
Participant 7: CounterQuill’s selective rewrite is helpful because I can keep what I wrote originally while also editing it part by part together with the AI … and the final response is created by CounterQuill and me together.
In addition to selective rewriting, many participants (12 — P2 - P6, P10, P11, P12, P17 - P20) reported that grammar checking was the first feature they would use. All participants (20) had thoughts on the empathetic rewriting feature: the empathetic rephrasing suggestions helped participants better express their ideas (17 — P1 - P12, P15 - P19), making their responses more relatable and humanizing (17 — P1 - P8, P10 - P13, P15 - P18, P20). The empathetic rephrasing suggestions helped participants better articulate their thoughts (11 — P2, P6 - P13, P19, P20), enhanced the impact of their counterspeech (15 — P1, P2, P3 - P9, P11 - P14, P18, P20) and give them emotional support (7 — P2, P3, P5, P7, P12, P14, P18). For example, P2 felt more connected and less confrontational in writing counterspeech and shared how the system improved the original counterspeech:
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=training, linewidth=0pt, backgroundcolor=training!10 {mdframed}
Participant 2: I think using this system to polish and refine my original draft really helped because it made my comment a little shorter but also made it more empathetic. It generally kept my content the same. The AI also added their own touch at the very end, something like ‘This is an opportunity for us to work together.’ I think that really added a sense of empathy and encouragement for mutual understanding. It doesn’t seem like you’re talking at them but talking with them.”
While some participants found the process of generating counterspeech with ChatGPT to require less effort (e.g., “ChatGPT can provide a quick response” - P14), they expressed a lack of ownership over the content generated by ChatGPT and were hesitant to post AI-generated counterspeech online (“I tried many prompts (with ChatGPT), but even if it gave me 10 different counterspeech, I wouldn’t use any of them … because they are too AI-like”). In contrast, participants (19 — P1-P19) preferred CounterQuill for future real-world online interactions. They particularly appreciated how CounterQuill incorporates tone adjustment while preserving their own voice and personal experiences (“I am likely to use CounterQuill in the future because its system features really keep my own word.” - P2).
7. DISCUSSION
Our findings provide insights for designing AI-assistant tools that empower users in online counterspeech writing. We discuss three key design implications in this context.
7.1. Encouraging User Ownership in AI-Assisted Writing
A major challenge in AI-mediated writing tools is maintaining a sense of personal ownership over AI-generated content (Louie et al., 2020; Mieczkowski, 2022). Although our study did not directly measure perceived ownership of the generated content, we found that participants’ satisfaction and willingness to post their written counterspeech online, which can be seen as a proxy of perceived ownership, were higher in CounterQuill’s condition. These findings echo prior research by Bhattacherjee et al. (2001), who found that higher levels of satisfaction create a sense of ownership of the results generated by the system, which in turn drives users to continue using the system (Bhattacherjee, 2001). When participants actively modified AI-generated content rather than simply accepting it, they reported a stronger sense of ownership over the counterspeech (§ 6.4.2). This was further reinforced when users incorporated personal experiences into their counterspeech and adjusted the AI’s suggestions to match their intended tone. Participants described the process as ”co-writing with the AI,” with some even referring to CounterQuill as a writing partner rather than just an automated tool.
Our work shows that integrating user-driven ideation and revision in AI-assisted counterspeech writing can increase users’ sense of ownership over their counterspeech. Prior research in HCI has shown that the more meaningful cognitive effort people invest in collaborating with AIMC systems, the stronger their sense of ownership towards the final output (Louie et al., 2021, 2020; Pierce et al., 2003; Mieczkowski, 2022). Our findings echo this: participants who invested effort in refining the AI’s suggestions—rather than passively accepting them—expressed a stronger satisfaction with their counterspeech. By designing AI assistance to prioritize user engagement in the writing process, AI-mediated communication (AIMC) tools like CounterQuill can empower users to take ownership of their responses rather than feeling as though they are merely transmitting AI-generated content.
Similarly, as highlighted in (§ 6.4.1), participants who had explicit control over CounterQuill’s customization settings were more likely to engage with and revise AI-generated suggestions. Research has shown that while AI-assisted writing can improve efficiency, excessive reliance on automation reduce user agency and engagement (Ping et al., 2024; Lee et al., 2022, 2024a). Our findings show that users value AI assistance that enhances their existing ideas rather than replacing them. To achieve this, AI writing assistants for counterspeech engagement should prioritize user-driven customization (Zhang et al., 2023), allowing users to selectively refine, rephrase, or adjust AI suggestions rather than accepting predefined outputs. Furthermore, embedding transparency mechanisms (e.g., visual indicators of AI contributions (Akhtar, 2024)) can help users retain control over their writing process, fostering trust and a greater sense of ownership (Mieczkowski, 2022; Draxler et al., 2024; Kim and Song, 2023).
7.2. Structuring AI Guidance to Reduce Cognitive Load
Effective counterspeech requires careful thought, yet users often experience cognitive overload when crafting responses to hate speech (Ping et al., 2024). CounterQuill’s step-by-step workflow—incorporating learning, brainstorming, and structured AI assistance—helps mitigate this challenge by breaking down complex tasks into manageable stages (§ 6.4). Echoing prior research (Ping et al., 2024), our findings suggest that AI writing assistants should avoid overwhelming users with unstructured suggestions; instead, they should provide guided interactions that progressively build toward the final output (Ping et al., 2024; Zhang et al., 2023).
Rather than providing fully generated responses, AI systems should support incremental co-creation, allowing users to shape and refine content instead of passively receiving suggestions (Caspi and Blau, 2011). Features like selective rewriting, interactive brainstorming, and context-aware feedback as demonstrated in CounterQuill help users integrate their own voice and perspective, increasing their willingness to engage in online discussions. By guiding users through the writing process, structured AI assistance can build confidence and encourage more deliberate, personalized counterspeech.
However, this structured guidance comes with a trade-off. CounterQuill’s multi-step workflow requires more time and effort than fully automated approaches like ChatGPT. While this level of engagement benefits users seeking to learn, brainstorm, and refine their responses (§ 6.4), it may not be suitable for experienced counterspeech writers who may want to prioritize speed to counteract and drown out online hate speech as quickly as possible before it spreads (Schieb and Preuss, 2016; ling Chung et al., 2023; Ping et al., 2024).
Our quantitative analysis (§6.1) showed that collaborating with ChatGPT required less cognitive effort and time than CounterQuill. However, in customized post-task questions, participants reported higher satisfaction with their final counterspeech when using CounterQuill, rating it 6.47 compared to ChatGPT’s 4.00—a 62% increase. Similarly, participants were 78% more likely to publish their counterspeech online when collaborating with CounterQuill (6.29 vs. 3.53). These findings suggest that while structured AI guidance demands more effort, it fosters deeper engagement and a stronger sense of ownership—key factors in promoting sustained participation in counterspeech efforts. Striking a balance between structured support and efficiency will be essential for designing AI systems that cater to both novice and experienced counter speakers.
7.3. Enhancing AI’s Role in Online Discourse Moderation
As platforms shift toward community-driven moderation (Brewer et al., 2020), AI-assisted counterspeech tools like CounterQuill can serve as more than just writing aids - they can actively shape how users engage with harmful content. Rather than replacing human content moderation with automated enforcement, AI can equip users with strategic communication skills to intervene in online discussions more effectively. This approach complements content moderation policies by fostering peer-led counter-messaging rather than relying solely on content removal.
Our findings indicate that structured AI guidance not only reduces cognitive load (§ 6.4.1) but also increases users’ confidence in their responses (§ 6.4.1), making them more willing to engage in public discourse. Participants appreciated AI-generated guidance that included contextual explanations alongside counterspeech suggestions, reinforcing their ability to engage in meaningful dialogue. Moreover, our results suggest that the greater the effort users invested in co-creating counterspeech, the more likely they were to share it publicly. This suggests that AI-mediated tools, when designed to prioritize user agency and strategic guidance, can encourage more active and constructive interventions in online discussions that may be fraught with hateful messages.
Finally, future AI-mediated moderation tools should focus not just on detecting and removing harmful content, but on educating and empowering users to respond constructively. By integrating adaptive AI guidance that responds to user expertise levels, these systems can help bridge the gap between passive content filtering and scalable, community-driven intervention strategies.
8. CONCLUSION
In this work, we introduced CounterQuill, an AI-mediated system designed to assist users in writing effective counterspeech. CounterQuill guides users through three key phases: a learning session to enhance understanding of hate speech and counterspeech, a brainstorming session to identify hateful elements and generate response strategies, and a co-writing session to refine counterspeech collaboratively with AI assistance.
Our user study demonstrated that CounterQuill effectively educates users about hate speech and empathy-driven counterspeech strategies. The brainstorming session, supported by AI feedback, increased users’ confidence in engaging with counterspeech, while the co-writing session enhanced their sense of self-efficacy and ownership. Notably, we found that the more effort users invested in collaborating with CounterQuill, the stronger their sense of ownership over the co-created content—leading to a higher likelihood of posting their responses on social media.
Limitations and future work. One limitation of our study is the potential influence of participants’ prior experience with AI and counterspeech on their engagement with CounterQuill. Additionally, our evaluation was limited to responses to a single hate speech example, which may not fully capture the system’s adaptability across different contexts. Future work should explore multiple hate speech scenarios to assess CounterQuill’s effectiveness more comprehensively.
Although participants indicated they would use CounterQuill to craft counterspeech, perceived effectiveness does not always translate into real-world adoption. However, the Theory of Planned Behavior (Ajzen, 1980, 1991) suggests that positive perceptions of a tool are strong predictors of future usage (Wan et al., 2014; Bosnjak et al., 2020). Given that users are unlikely to adopt a tool they perceive as ineffective, our findings indicate that CounterQuill holds promise for real-world impact.
For future research, we propose developing CounterQuill as a browser extension (e.g., for Chrome and Edge) and integrating it into social media platforms like Reddit and X. Additionally, we plan to conduct a longitudinal study to evaluate CounterQuill extension’s usability and effectiveness in real-world online discourse.
Appendix A Preliminary survey questions
  1. (1)
    Have you used Large Language Model (LLM) tools or an AI chatbot, such as ChatGPT, Google Gemini or Claude? (Yes, No)
  2. (2)
    You have indicated that you have used an AI chatbot, such as ChatGPT. Can you briefly describe this experience?
  3. (3)
    Have you ever countered hate content on online social media? (Yes, No)
  4. (4)
    You have indicated that you have countered hate speech on social media. Can you briefly describe this experience?
  5. (5)
    If you were writing a counterspeech on social media, would you use AI technology like ChatGPT to help you? (1 = Definitely not, 5 = Definitely yes)
  6. (6)
    Why would or wouldn’t you use AI technology to help you write a counterspeech?
  7. (7)
    What is your age?
  8. (8)
    What is your gender identity? Select as many that apply.
  9. (9)
    What is your ethnicity?
  10. (10)
    What is your education level?
  11. (11)
    What is your sexual orientation?
  12. (12)
    In general, how would you describe your political views?
Appendix B Guiding interview questions
  1. (1)
    How do you perceive implicit hate speech after reviewing the learning materials provided?
  2. (2)
    What are your thoughts on the negative impacts of explicit versus implicit hate speech?
  3. (3)
    How would you describe empathy-based counterspeech, and what role do you think it plays in addressing hate speech?
  4. (4)
    What were your impressions of using colors to highlight targeted identities and dehumanizing actions? How did this influence your understanding?
  5. (5)
    Do you feel that your ability to identify targeted identities and dehumanizing language has evolved through this exercise? If so, in what ways?
  6. (6)
    How did the brainstorming questions influence your approach to responding to hate speech compared to your initial reaction?
  7. (7)
    How did you find the personalized feedback provided by the system? In what ways did it help or hinder your process?
  8. (8)
    How did you perceive the AI’s ”polish” and ”rewrite as empathetic” features in refining your response?
  9. (9)
    How do you value the ability to selectively rewrite specific sentences using AI compared to other editing tools?
  10. (10)
    How did the combination of learning, brainstorming, and writing aids affect your confidence in crafting counterspeech?
Appendix C Post-Study NASA Question
To save space on the page, we have combined the post-study questions, as the content of the questions is identical, with the only difference being that Condition A refers to GPT and Condition B refers to CounterQuill.
Please rate your experience using ChatGPT (condition A) / CounterQuill (condition B) to write counter-speech on the following dimensions. For each dimension, select a number from 1 (very low) to 7 (very high).
  1. (1)
    Please enter your participation number.
  2. (2)
    How much mental activity was required to use ChatGPT (condition A) / CounterQuill (condition B) for counter-speech writing (e.g., how much thinking or deciding was required to perform the task)?
  3. (3)
    How much physical activity was required to use ChatGPT (condition A) / CounterQuill (condition B) for counter-speech writing (e.g., typing, clicking, etc.)?
  4. (4)
    How much time pressure did you feel due to the pace at which the task of counter-speech writing using ChatGPT (condition A) / CounterQuill (condition B) occurred?
  5. (5)
    How hard did you have to work (mentally and physically) to accomplish your level of performance in using ChatGPT (condition A) / CounterQuill (condition B) for counter-speech writing?
  6. (6)
    How successful do you think you were in accomplishing the goals of the task (writing effective counter-speech using ChatGPT (condition A) / CounterQuill (condition B))?
  7. (7)
    How insecure, discouraged, irritated, stressed, and annoyed versus secure, gratified, content, relaxed, and complacent did you feel during the task of using ChatGPT (condition A) / CounterQuill (condition B) for counter-speech writing?
Appendix D Post-Study Custom Question
Please rate your experience using ChatGPT (condition A) / CounterQuill (condition B) to write counter-speech on the following dimensions. For each question, select a number from 1 (strongly disagree) to 7 (strongly agree).
  1. (1)
    How confident are you in your ability to identify online hate speech that dehumanizes individuals or groups based on their identity after co-writing with ChatGPT (condition A) / CounterQuill (condition B)?
  2. (2)
    How effective is ChatGPT (condition A) / CounterQuill (condition B) at helping you with step-by-step guides for planning and thinking about counter-speech writing?
  3. (3)
    To what extent did using ChatGPT (condition A) / CounterQuill (condition B) increase your self-efficacy (belief in your own abilities) in writing effective counter-speech?
  4. (4)
    How engaged and motivated did you feel while using ChatGPT (condition A) / CounterQuill (condition B) to write counter-speech?
  5. (5)
    How satisfied are you with the final counter-speech you created using ChatGPT (condition A) / CounterQuill (condition B)?
  6. (6)
    If you had the opportunity, how likely would you be to post this counter-speech created using ChatGPT (condition A) / CounterQuill (condition B) on social media?
Appendix E Quiz questions
  • Q1.
    Which statement accurately reflects the nature of hate speech?
    • A.
      Hate speech primarily targets a person’s actions or choices.
    • B.
      Hate speech is any form of speech that someone finds offensive.
    • C.
      Hate speech specifically targets core aspects of a person’s or group’s inherent identity, such as race or religion.
    • D.
      Hate speech is the same as constructive criticism but with a negative tone.
  • Q2.
    Which of the following is NOT a consequence of implicit hate speech?
    • A.
      Desensitization to derogatory language, resulting in reduced empathy for victims and insensitivity towards hate speech.
    • B.
      Promoting important and constructive diverse viewpoints, fostering understanding and tolerance among different groups.
    • C.
      Inducing psychological and emotional harm, increasing feelings of fear, anger, and helplessness for individuals targeted by implicit hate speech.
    • D.
      Promoting prejudice and fostering bigotry and intolerance through stereotyping and negative perceptions of certain groups.
  • Q3.
    Which statement does not accurately reflect the role of counterspeech?
    • A.
      Counterspeech can be risky (e.g., leading to backlash).
    • B.
      Counterspeech can help people to change their hateful perspectives. .
    • C.
      Counterspeech can encourage others to also counter hate.
    • D.
      Counterspeech can always exacerbate tensions and escalate the conflict.
  • Q4.
    Which of the following statements about empathy-based counterspeech aligns with the given learning materials?
    • A.
      Empathy-based counterspeech is most effective when it involves retaliatory attacks against the attacker.
    • B.
      Empathy-based counterspeech humanizes the victim and highlights the potential harm caused by the attacker’s behavior.
    • C.
      Empathy-based counterspeech is ineffective unless it includes aggressive language and strong condemnations.
    • D.
      Empathy-based counterspeech is only effective when delivered by a single user rather than a collective group.
Appendix F CounterQuill prompt
Because CounterQuill prompts are long, we created a GitHub repository to save our prompts. See https://github.com/AnonymousUseOnly/CounterQuill_Prompt for details.
Appendix G Results of Welch’s Two-Sample t-tests
We conducted Welch’s Two-Sample t-tests that accounts for unequal variances between groups (Sakai, 2016), as detailed in Tables A1 and A2. The results confirmed that the sequence—whether participants began with ChatGPT or CounterQuill—did not significantly influence their performance or feedback on the system.
Table A1. Welch’s Two-Sample t-tests comparing the effect of system usage order (ChatGPT First vs. CounterQuill First) on NASA-TLX workload dimensions (Mental, Physical, Temporal Demands, Performance, Effort, and Frustration). No significant differences were found (all p ¿ 0.05), indicating order did not affect participants’ performance or feedback.
Platform NASA-TLX Items $Mean\pm SD$ t-value p-value 95% CI
ChatGPT First CounterQuill First
ChatGPT Mental Demand 4.80$\pm$1.32 4.30$\pm$1.95 0.673 0.511 [-1.077, 2.077]
Physical Demand 2.40$\pm$1.65 1.90$\pm$1.52 0.705 0.490 [-0.991, 1.991]
Temporal Demand 2.90$\pm$1.37 2.30$\pm$1.49 0.936 0.362 [-0.748, 1.948]
Performance 4.00$\pm$1.05 3.80$\pm$1.14 0.408 0.688 [-0.830, 1.230]
Effort 2.60$\pm$1.17 2.20$\pm$0.92 0.849 0.408 [-0.594, 1.394]
Frustration 3.40$\pm$1.90 2.60$\pm$2.17 0.878 0.392 [-1.118, 2.718]
CounterQuill Mental Demand 5.50$\pm$0.53 4.50$\pm$1.65 1.826 0.096 [-0.208, 2.208]
Physical Demand 3.90$\pm$1.66 2.80$\pm$1.69 1.468 0.159 [-0.474, 2.674]
Temporal Demand 4.90$\pm$1.73 4.10$\pm$1.20 1.203 0.246 [-0.610, 2.210]
Performance 6.20$\pm$0.79 6.50$\pm$0.71 -0.896 0.382 [-1.004, 0.404]
Effort 4.20$\pm$1.55 3.90$\pm$1.79 0.400 0.694 [-1.276, 1.876]
Frustration 1.70$\pm$0.82 1.30$\pm$0.67 1.188 0.251 [-0.309, 1.109]
Table A2. Welch’s Two-Sample t-tests comparing the effect of system usage order (ChatGPT First vs. CounterQuill First) on responses to customized questions measuring HS Identification Confidence, Brainstorming Effectiveness, Self-Efficacy in CS Writing, Engagement with AI, Satisfaction with Counterspeech, and Willingness to Post Online. No significant differences were found (all p ¿ 0.05), indicating order did not influence participants’ responses.
Platform Category of Question $Mean\pm SD$ $t$-value $p$-value 95% CI
ChatGPT First CounterQuill First
ChatGPT HS Identification Confidence 5.00$\pm$1.41 3.40$\pm$1.96 2.097 0.052 [-0.014, 3.214]
Brainstorming Effectiveness 3.50$\pm$1.27 3.00$\pm$1.49 0.808 0.430 [-0.803, 1.803]
Self-Efficacy in CS Writing 4.00$\pm$1.76 2.60$\pm$1.65 1.835 0.083 [-0.204, 3.004]
Engagement with AI 4.10$\pm$1.52 3.00$\pm$1.70 1.524 0.145 [-0.418, 2.618]
Satisfaction with CS 4.50$\pm$1.43 3.30$\pm$1.42 1.882 0.076 [-0.140, 2.540]
Willingness to Post Online 4.10$\pm$1.60 2.80$\pm$1.14 2.100 0.052 [-0.011, 2.611]
CounterQuill HS Identification Confidence 6.40$\pm$0.70 6.20$\pm$0.79 0.600 0.556 [-0.501, 0.901]
Brainstorming Effectiveness 6.60$\pm$0.52 6.20$\pm$0.79 1.342 0.199 [-0.234, 1.034]
Self-Efficacy in CS Writing 6.40$\pm$0.97 6.10$\pm$0.57 0.847 0.411 [-0.457, 1.057]
Engagement with AI 6.00$\pm$0.94 5.90$\pm$0.74 0.264 0.795 [-0.699, 0.899]
Satisfaction with CS 6.60$\pm$0.52 6.30$\pm$0.48 1.342 0.196 [-0.170, 0.770]
Willingness to Post Online 6.30$\pm$0.82 6.20$\pm$0.79 0.277 0.785 [-0.658, 0.858]
Appendix H Examples of hate speech
Warning: Table LABEL:speech_examples_table contains potentially offensive or upsetting hate speech examples used in the user study. These are presented for academic and research purposes only and do not reflect the author’s views.
Table A3. Examples of Hate Speech and Associated Topics
Topic Example
#gender Women are naturally better at caring for others and doing housework than men, who are a mess and illogical at work.
#disability ”Mentally disabled people should not be allowed to vote. They are a burden to society.”
#race ”I saw a random black man jogging nearby our house today and now I feel unsafe walking around my own neighborhood.”
#sexual_orient Lesbians are often stereotyped as ’tomboys,’ rejecting traditional female roles and appearances, and when I see them, I get uncomfortable because they’re not feminine at all.
#race I guessed you must be Japanese because in my experience Japanese people have very modern haircuts and clothing. Chinese people usually have very traditional looking clothing and hair, and you don’t look that at all! You look so chic and modern!
#gender ”It’s impressive how you handle such a demanding job and still manage to keep your family together. I guess it’s rare to find women who can balance both without letting one slip.”
#race ”Indeed, you are doing fine in a country with a majority white population. How is Mexico faring these days? People come here because Mexico couldn’t develop properly.”
#race It’s amazing how those kids from the inner city manage to graduate high school. I mean, considering their circumstances, you’d expect them not to achieve much.
#disability ”Persons with learning difficulties should be discouraged to attend primary school and they should be excluded from education.”
#sexual_orient ”Stop tucking your hair into your hat, you’ll attract a lot of lesbians that will want to be your girlfriend.”
#gender ”Women are often seen as being too emotional to handle leadership roles effectively.”
#race ”I’ve noticed that neighborhoods tend to go downhill when more of black people move in, you know? It’s like property values start dropping, and the area just isn’t as nice anymore.”
#disability ”Children with Down Syndrome should be isolated, because in any case normal children would make them feel different, ending up in unwanted discrimination.”
#sexual_orient ”Gays are sick. They should be isolated and isolated soon. They are a danger to society.”
#religion ”Islamists/Muslims have mastered the corrupt art of playing the victim and blaming other cultures for the violent beliefs taught in their main books.”
#religion ”Right. I’m questioning if this was a legitimate event, though. It looks like a hoax perpetrated by the media, which is perceived to be owned predominantly by Jewish interests, to be honest.”
#religion ”Every time I board a plane and see someone wearing a turban, I can’t help but feel a bit nervous, as if the safety of the flight is suddenly in question. It’s just an instinctive reaction, I guess.”
#race ”It was cleaner when white people lived here. Then the Blacks came in and now look! There’s trash all over the streets, dogs running loose, and the smell of smoke every sunny day.”